### v1

# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

Abstract
We consider repeated allocation of a shared resource via a non-monetary mechanism, wherein a single item must be allocated to one of multiple agents in each round. We assume that each agent has i.i.d. values for the item across rounds, and additive utilities. Past work on this problem has proposed mechanisms where agents can get one of two kinds of guarantees: $(i)$ (approximate) Bayes-Nash equilibria via linkage-based mechanisms which need extensive knowledge of the value distributions, and $(ii)$ simple distribution-agnostic mechanisms with robust utility guarantees for each individual agent, which are worse than the Nash outcome, but hold irrespective of how others behave (including possibly collusive behavior). Recent work has hinted at barriers to achieving both simultaneously. Our work however establishes this is not the case, by proposing the first mechanism in which each agent has a natural strategy that is both a Bayes-Nash equilibrium and also comes with strong robust guarantees for individual agent utilities.
Our mechanism comes out of a surprising connection between the online shared resource allocation problem and implementation theory. In particular, we show that establishing robust equilibria in this setting reduces to showing that a particular subset of the Border polytope is non-empty. We establish this via a novel joint Schur-convexity argument. This strengthening of Border’s criterion for obtaining a stronger conclusion is of independent technical interest, as it may prove useful in other settings.
1 Introduction
Consider a single indivisible public resource being repeatedly allocated between multiple agents – for example, a scientific instrument shared by multiple university labs. Since the resource is public, its allocation should be determined ideally without using monetary transfers. Moreover, the principal wants to allocate the resource in a way that is both efficient (so labs get the instrument only when they have great need for it) and also fair (so that the amount of time each lab gets to use it is roughly proportional to some pre-determined share). Each individual agent is of course self-serving, and so some mechanism is required to encourage agents to request for the resource only when they need it the most. And in the absence of money, the principal may not be able to verify agents’ reports, or have any strong beliefs about their valuations and actions.
The above problem was first studied under a simple model by [guo2009competitive], with a single indivisible item per round, and agents with random valuations across rounds. Early work on this setting [guo2009competitive, cavallo2014incentive, balseiro2019multiagent, gorokh2021monetary] focused on the question of how efficient a non-monetary mechanism could be in such a setting; as we discuss in Section 1.3, this led to several mechanisms with near-efficient Bayes-Nash equilibria. All these mechanisms, however, critically depend on knowing the exact value distributions.
[…]
So is it possible to design mechanisms that unite the above two streams for the setting of [guo2009competitive]? We define this formally in Section 2.2, but at a high level, we seek mechanisms that support a good robust equilibrium: a strategy that simultaneously satisfies the following:
1. Equilibrium Performance: Assuming all agents follow the strategy, no one has an incentive to deviate. Moreover, under this equilibrium, every agent enjoys high utility. 2. Robust Performance: If an agent follows the strategy, but others act arbitrarily (maybe even collusively), then the agent still enjoys some (relatively high) utility guarantee.
This definition is of course under-specified in terms of what we mean by ‘good’: note that never allocating, or allocating uniformly at random, supports any strategy both as an equilibrium, and also gives every agent a (weak) robust guarantee. Our main result however is a new mechanism for repeated non-monetary allocation which admits a simple strategy that is simultaneously a Bayes-Nash equilibrium in the infinite-horizon limit (and close to Bayes-Nash equilibrium in the finite case), and also matches the current best robustness guarantee. Moreover, we achieve this via a surprising strengthening of Border’s theorem, which may prove useful in other settings.
1.1 Overview of our Mechanism and Main Result
[…]
Our aim is to develop a mechanism that admits a robust equilibrium: a joint profile of simple strategies for the agents $(\pi_{1},\pi_{2},\ldots,\pi_{n})$ that is simultaneously a Bayes-Nash equilibrium where every player receives per-round utility at least $\lambda_{\mathtt{NASH}}v^{\star}_{i}$ (which we refer to as a $\lambda_{\mathtt{NASH}}$-good Bayes-Nash equilibrium) and at the same time the same strategy ${\pi}_{i}$ guarantees that agent $i$ receives per-round utility at least $\lambda_{\mathtt{ROB}}v^{\star}_{i}$, irrespective of how other agents act, including possibly collusive actions (which we call $\lambda_{\mathtt{ROB}}$-robust). We want strategies to be simple so that they are more meaningful in practice. Note also that since we benchmark against the ideal utility $v_{i}^{\star}$ instead of the first-best outcome, we may not be able to get $\lambda_{\mathtt{NASH}}=1$. For example, in the case where all agents have equal shares and i.i.d values $V_{i}[t]\sim\textrm{Bernoulli}(1/n)$, the best possible is $\lambda_{\mathtt{NASH}}\leq 1-(1-1/n)^{n}\approx 1-1/e$ [banerjee2023robust].
Within this backdrop, we propose a new mechanism we call Budgeted Robust Border (BRB) (see Mechanism 1). At a high level, the mechanism has two simple components:
  1. 1.
    Budget-regulated bidding: Each agent has a budget of $B_{i}\approx\alpha_{i}T+o(T)$ tokens, which regulates the number of times she can bid for the item over $T$ rounds.
  2. 2.
    Probabilistic Allocation: At time $t$, given the set of bidders $S[t]=S\subseteq[n]$, the principal allocates the item to some agent $i\in S$ according to some pre-decided probability $p_{i}^{S}$.
For the exact expression for the budget $B_{i}$ see Proposition 4.1; the additional $o(T)$ budget is needed to get high probability guarantees. The allocation probabilities $\{p_{i}^{S}\}$ are more complex (and indeed, that is where the main novelty of our work lies, as we discuss next). However, at this point, we can already summarize our main result, as follows (see Theorem 4.3):
Under the BRB Mechanism, the policy where player $i$ bids in round $t$ whenever her value $V_{i}[t]$ is in the top $\alpha_{i}$-quantile (and subject to her budget), is simultaneously: Robust with factor $\lambda_{\mathtt{ROB}}\geq 1/2$ A Bayes-Nash equilibrium with $\lambda_{\mathtt{NASH}}=1-\prod_{j=1}^{n}(1-\alpha_{j})\geq 1-1/e$
In Lemma 2.1 we prove the above $\lambda_{\mathtt{NASH}}$ is minimax optimal for any vector of fair shares $\{\alpha_{i}\}_{[n]}$; $\lambda_{\mathtt{ROB}}=1/2$ matches the current best [gorokh2019remarkable, banerjee2023robust, fikioris2023online].
1.2 Overview of our Techniques
Our main conceptual idea is a connection between repeated non-monetary allocation and implementation theory (in particular, Border’s theorem). Our main technical novelty is a way to modify the Border flow network in order to select $1/2$-robust equilibria. Our mechanism uses three main ideas: $(i)$ a repeated all-pay mechanism to control bidding rates, $(ii)$ guaranteeing good interim allocations via the regular Border condition, and $(iii)$ strengthening the Border condition to guarantee robustness. We now describe these in brief.
We start by first trying to get good equilibria. In the infinite horizon setting, our idea is as follows: We restrict agents to only bid or not in each round, and further restrict them to bid at most an $\alpha_{i}$ fraction of the time. Our mechanism aims to guarantee agent $i$ a good interim allocation conditioned on them bidding (the probability that they are allocated), assuming every other agent $j$ bids independently with probability at most $\alpha_{j}$. Finally, to pass to the finite horizon, we use an idea from [gorokh2021monetary] and employ an all-pay mechanism with budgets $\{\alpha_{i}T+o(T)\}$, to ensure that agents are not incentivized to bid at a rate higher than $\alpha_{i}$, while also not running out of budget with high probability.
To ensure good interim allocations, we use probabilistic allocation with preset probabilities $p_{i}^{S}$ for assigning the item to agent $i$ when the set $S$ of agents bid for each subset $S\subseteq[n]$, and agents $i\in S$. These are chosen to ensure each agent’s bids are accepted with probability at least $\lambda_{\mathtt{NASH}}$, assuming each agent $i$ bids independently with probability $\alpha_{i}$. The bid budgets and probabilistic allocation serves to somewhat insulate each agent from the actions of others. Consequently, in equilibrium, agent $i$’s best response is to bid when her value is in the highest $\alpha_{i}$ quantile, and which makes bids independent across agents (since valuations are independent). Border’s theorem (see Theorem 5.2) provides necessary and sufficient conditions for which interim allocation probabilities are feasible; in particular, we use this to show we can always guarantee $\lambda_{\mathtt{NASH}}\geq 1-\prod_{j}(1-\alpha_{j})\geq(1-1/e)\approx 0.63.$ (see Theorems 4.3 and 4.2). Our guarantee is in fact minimax optimal for any vector of fair shares $\{\alpha_{i}\}_{[n]}$: in Lemma 2.1 we show a valuation profile that gives a matching upper bound even if the principal knows agents’ realized values.
The challenge however is that many interim allocation probabilities admitted by Border’s theorem are not robust: if all other agents collude against agent $i$, they can severely limit agent $i$’s allocation. In fact, different allocation rules inducing the same interim allocation can have vastly different robustness guarantees (see Section 6.2). Even computing $\lambda_{\mathtt{ROB}}$ for a given allocation rule appears difficult, as we need to search over all potentially correlated bidding schemes by the other agents.
Our main technical novelty is in identifying the subset of interim allocations allowed by Border’s theorem which are $1/2$-robust. To do this, we consider any chosen agent $i$, and characterize the bang-per-buck (in terms of blocking agent $i$ per bid token spent) of each possible set $S\subseteq[n]\setminus\{i\}$ of colluding agents bidding in a round. Using this, we show that the critical case is allocating when only 2 agents bid (when agents collude, they can make this the dominant case). We show that it is possible to make the mechanism $1/2$-robust (matching the current best robustness guarantee), by modifying the Border flow network by adding appropriate bounds to edges going out of each doubleton subset $\{i,j\}\subset[n]$ (Fig. 2). This leads to a modification of Border’s criterion for ensuring the resultant network still supports the same maximum amount of flow (Lemma 6.2). Finally, in Lemma 6.3, we demonstrate the resultant polytope is non-empty, via establishing the Schur-convexity/concavity of a certain bivariate function that comes out of our new criterion.
Taken together, our arguments demonstrate allocation probabilities that simultaneously guarantee for all agents at least an $(1-1/e)$ fraction of their ideal utility at Nash equilibrium and, at the same time, a $1/2$ fraction of their ideal utility even if other agents collude. While the allocation rules in the general case are somewhat involved, we illustrate them explicitly in two simple corner cases – $2$ agents with arbitrary shares in Section 3, and $n$ agents with equal shares in Section 6.
1.3 Related Work
[…]
2 Preliminaries
2.1 Model
We consider the following simple setting of repeatedly allocating a single, indivisible item among $n$ agents; this was first introduced in the work of [guo2009competitive], and has since been widely studied, as we discuss in Section 1.3. At each discrete time-step $t=1,2,\dots,T$, a principal receives a new item and must select an agent $i\in[n]$ to allocate the item to (or not allocate the item).
We assume that each agent $i$ has a private value $V_{i}[t]$ for the item at time $t$, where $V_{i}[t]\sim\mathcal{F}_{i}$ for some value distribution $\mathcal{F}_{i}$ not depending on $t$. We assume that the $V_{i}[t]$ are nonnegative, bounded, and independent across both agents and time. Let $X_{i}[t]$ be the indicator that agent $i$ was allocated the item at time $t$. At time $t$, the agent gets utility $U_{i}[t]=V_{i}[t]X_{i}[t]$. We assume that utilities are additive over time so that an agent’s total utility after $T$ time periods is $\sum_{t=1}^{T}U_{i}[t]$.
2.2 Ideal Utility and Benchmarks
As in previous work in this setting [gorokh2019remarkable, banerjee2023robust, fikioris2023online], we assume each agent $i$ has an exogenously defined fair share $\alpha_{i}$, where each $\alpha_{i}$ is nonnegative and $\sum_{i=1}^{n}\alpha_{i}=1$. The fair share $\alpha_{i}$ of agent $i$ represents the fraction of rounds we want agent $i$ to win the item under ideal circumstances. As done by previous work, we use the benchmark of ideal utility. The ideal utility of an agent $i$ is the maximum expected utility they can obtain from a single round if they can obtain the item simply by requesting it, but they are only allowed to request it with probability at most their fair share $\alpha_{i}$. Formally, the (per-round) ideal utility is the following.
Definition 2.1 (Ideal Utility).
Agent $i$’s ideal utility is the value of the following maximization problem over measurable $\rho_{i}:[0,\infty)\to[0,1]$:
$\begin{split}v_{i}^{\star}=\max\,\operatorname*{\mathbb{E}}_{V_{i}\sim\mathcal{F}_{i}}[V_{i}\rho_{i}(V_{i})]\quad\text{subject to}\quad\operatorname*{\mathbb{E}}_{V_{i}\sim\mathcal{F}_{i}}[\rho_{i}(V_{i})]\leq\alpha_{i}.\end{split}$
[…]
We study what fraction of her ideal utility an agent $i$ can guarantee in two settings: First, we examine the performance of Bayes-Nash equilibria, where each agent is trying to maximize their own utility. Second, we examine what fraction of her ideal utility agent $i$ can guarantee robustly, i.e., even if the other agents $j\neq i$ are playing adversarially and collude to harm agent $i$ without regard for their own utilities. Critically, we want an agent to achieve both the above with the same (ideally simple) strategy. As in Section 1, we want a mechanism that achieves the following.
$\lambda_{\mathtt{NASH}}$-Nash $\lambda_{\mathtt{ROB}}$-robust equilibria: A mechanism with policy profile $(\pi_{1},\pi_{2},\ldots,\pi_{n})$ s.t. $(\pi_{1},\pi_{2},\ldots,\pi_{n})$ forms a Nash equilibrium as $T\to\infty$ in which each agent $i$ gets a $\lambda_{\mathtt{NASH}}$ fraction of their ideal utility. Policy $\pi_{i}$ is $\lambda_{\mathtt{ROB}}$-robust with $\lambda_{\mathtt{ROB}}$ is not much smaller than $\lambda_{\mathtt{NASH}}$: it guarantees agent $i$ a $\lambda_{\mathtt{ROB}}$ fraction of her ideal utility even if agents $j\neq i$ collude and act adversarially.
To understand how the ideal utility benchmark behaves under Bayes-Nash equilibria, we first present a minimax bound on what fraction $\lambda_{\mathtt{NASH}}$ of ideal utility agents can get under ideal conditions. This will help later certify the minimax optimality of our mechanism in the equilibrium setting. Note that this bound is mechanism-agnostic, as in our example below, the mechanism has access to the agents’ realized values.
Lemma 2.1.
If each agent $i$ has value distribution $\mathcal{F}_{i}=\mathrm{Bernoulli}(\alpha_{i})$, it is impossible to guarantee each agent $i$ a $\lambda$ fraction of their ideal utility in expectation for $\lambda>1-\prod_{j\in[n]}(1-\alpha_{j})$, even if the mechanism knows $V_{i}[t]$ for all $i,t$ before round $1$. Note that $\inf_{n,\alpha_{1},\ldots,\alpha_{n}}1-\prod_{j\in[n]}(1-\alpha_{j})=1-1/e$.
Proof.
Under these Bernoulli value distributions, the probability that some agent has value $1$ for the item in a given round is $1-\prod_{j\in[n]}(1-\alpha_{j})$, so under any allocation rule,
$\frac{1}{T}\sum_{t=1}^{T}\sum_{i\in[n]}\mathbb{E}[U_{i}[t]]\leq 1-\prod_{j\in[n]}(1-\alpha_{j}).$ (1)
With these Bernoulli value distributions, each agent’s ideal utility is exactly $\alpha_{i}$. If each agent is obtaining a $\lambda$ fraction of their ideal utility in expectation, then
$\frac{1}{T}\sum_{t=1}^{T}\sum_{i\in[n]}\mathbb{E}[U_{i}[t]]\geq\lambda\sum_{i\in[n]}\alpha_{i}=\lambda,$
so $\lambda\leq 1-\prod_{j\in[n]}(1-\alpha_{j})$ by (1). ∎
3 Warm Up: Repeated Allocation with Two Agents
To motivate the design of our mechanism and show how we can achieve good robust equilibria as outlined in Section 2.2, we first give a simplified version of our mechanism in the special case where there are only two agents. Our mechanism lets each agent decide to bid for the resource or not in each round, up to some limited number of bids. It makes sense that when only one agent bids (and has not exceeded her bid budget), she is guaranteed the item. What is unclear is who gets allocated when both agents bid. We show that carefully randomizing this allocation leads to the optimal guarantee of Lemma 2.1, both in terms of the equilibrium and the robust setting, i.e., $\lambda_{\mathtt{NASH}}=\lambda_{\mathtt{ROB}}=1-(1-\alpha_{1})(1-\alpha_{2})$. To extend this result to $n$ agents, we need to set exponentially many parameters (the allocations when any subset $S$ agents bids), which we do by strengthening Border’s theorem in Section 5.
Our mechanism proceeds as follows: Each agent $i$ starts with $\alpha_{i}T$ bid tokens. At each round, each agent can bid 00 or $1$. If both agents bid 00, no one receives the item. If only one agent bids $1$, then she receives the item in that round. If both agents bid $1$, then we allocate the item to agent $1$ with probability $p_{1}^{\{1,2\}}=p$ and to agent $2$ with probability $p_{2}^{\{1,2\}}=1-p$. Regardless of who wins the item, each agent $i$ pays their bid. If an agent’s budget becomes 00, they can not bid.
Suppose agent $1$ has a fair share $\alpha_{1}=\alpha$ and agent $2$ has a fair share $\alpha_{2}=1-\alpha$. For simplicity, we assume that agent $i$ has value distribution with a continuous CDF $F_{i}$, and assume that the budget constraint is enforced only in expectation. (This is to help simplify the description of this warm-up case; we later fix this using inflated budgets, as in [gorokh2021monetary]). We now show that there is a simple Nash equilibrium for any $p$, where each agent bids only when their value is in the top $\alpha_{i}$-quantile of her value distribution.
Proposition 3.1 (Informal).
Each agent $i$ bidding at time $t$ if her value $V_{i}[t]$ is in the top $\alpha_{i}$-quantile of her value distribution is a Nash equilibrium.
Proof sketch.
First, observe that the expected spending of each agent $i$ under this strategy profile is indeed at most $B_{i}[1]=\alpha_{i}T$, since each agent will be bidding i.i.d. $\mathrm{Bernoulli}(\alpha_{i})$. Next, note that there is nothing that agent $1$ can do to affect player $2$’s behavior. She must solve her own stochastic control problem where her policy does not affect agent $2$’s behavior. At any time $t$, if agent $1$ bids, her probability of winning is $p_{1}=\alpha+(1-\alpha)p$, where the $\alpha$ term is the probability that agent $2$ does not bid, and the $(1-\alpha)p$ term is the probability that agent $2$ bid, and agent $1$ wins give both agents bid. Thus, agent $1$’s control problem reduces to selecting $\alpha_{1}T$ time periods in which to bid, and she wins a $p_{1}$ fraction of them in expectation. Agent $1$ should bid on her highest-valued $\alpha_{i}T$ time periods, which is precisely what bidding whenever her value $V_{i}[t]$ is in the top $\alpha_{i}$-quantile of her distribution does (in expectation). ∎
We next pick $p$. Note that agent $1$’s expected utility in the above equilibrium is
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{1}[t]]=\frac{1}{T}\sum_{t=1}^{T}p_{1}\mathbb{E}[V_{1}[t]\boldsymbol{1}\{V_{1}>F_{1}^{-1}(1-\alpha_{1})\}]=p_{1}v_{1}^{\star}$
where $p_{1}=\alpha+(1-\alpha)p$ is the probability agent $1$ wins a round conditioned on her bidding. Symmetrically, player $2$ gets expected utility $\frac{1}{T}\sum_{t}\mathbb{E}[U_{2}[t]]=p_{2}v_{2}^{\star}$ where $p_{2}=(1-\alpha)+\alpha(1-p)=1-\alpha p$. Setting $p=1-\alpha$ guarantees both agents $1-\alpha(1-\alpha)$ fraction of their ideal utility. This achieves $\lambda_{\mathtt{NASH}}=1-\alpha(1-\alpha)$ and by Lemma 2.1, this is the best factor any allocation rule can get.
The choice of $p=1-\alpha$ might seem unusual, especially when $\alpha$ is close to 00 or $1$ where one agent wins most times. The intuition behind this is as follows: if agent $1$ has small fair share and agent $2$ has a high one, then using a high $p$ and favoring agent $1$ has a small effect on agent $2$, most of whose bids are uncontested. Therefore, if we want equal outcomes, i.e., $p_{1}=p_{2}$, we have to favor the agent with the small fair share, breaking most ties in her favor.
In this special case, it is not hard to argue that agent $1$ enjoys the same utility guarantee even if agent $2$ acts adversarially. Since agent $1$’s bids are i.i.d. across time, the worst agent $2$ can do is bid as much as her budget constraint allows, which is the same as her equilibrium behavior. Using a symmetric argument for agent $2$, we get the following proposition.
Proposition 3.2 (Informal).
In the above mechanism with $p=1-\alpha$, agent $i$ bidding when her value is in her top $\alpha_{i}$-quantile is a $\lambda_{\mathtt{ROB}}$-robust strategy with $\lambda_{\mathtt{ROB}}=\quantity\big(1-\alpha_{i}(1-\alpha_{i}))$.
This means we achieve $\lambda_{\mathtt{ROB}}=1-(1-\alpha_{1})(1-\alpha_{2})$ which also matches the upper bound of any allocation of Lemma 2.1.
In coming sections, we generalize this mechanism for $n\geq 2$ players, where the main difficulty is resolving concurrent bids for any subset $S\subseteq[n]$ with $|S|\geq 2$. Since there are exponentially many such subsets, it is much harder to find a good or simple allocation rule that has the same properties as our choice of $p$ above. We will show that we can get the same optimal factor under equilibrium performance, $\lambda_{\mathtt{NASH}}=1-\prod_{i}(1-\alpha_{i})$, using Border’s theorem. For the robust guarantee, we show that a simple application of Border’s theorem is not enough. Instead, we have to consider more restrictions on how we handle concurrent bids, which leads to the strengthening of Border’s theorem in Lemma 6.2. Unfortunately, the robust guarantee weakens for $n>2$, and we get $\lambda_{\mathtt{ROB}}\approx 1/2$, which we show is tight under any rule for resolving concurrent bids.
4 Mechanism and Proposed Strategy
Input: Fair shares $\alpha_{i}$, number of rounds $T$, allocation probabilities $(p_{i}^{S})_{i\in S},\forall S\subseteq[n]$, budget slack $\delta_{i}^{T}$.
Endow each agent with a budget $B_{i}[1]=\alpha_{i}(1+\delta_{i}^{T})T$ of bid tokens;
for $t=1,2,\dots,T$ do
      Agents submit bids $b_{i}^{t}\in\{0,1\}$;
      Budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $B_{i}[t]\leq 0$;
      Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents;
      A winner $i^{t}$ is randomly selected from $S[t]$ according to a probability distribution $(p_{i}^{S[t]})_{i\in S[t]}$;
      Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$;
      end for
MECHANISM 1 Budgeted Robust Border (BRB) Mechanism
We now extend our mechanism from Section 3 and present it for the general case of $n\geq 2$. Similar to Section 3, every agent $i$ bids for the item or not in every round and can make (approximately) at most $\alpha_{i}T$ bids in total. In every round where agents bid for the resource, our mechanism randomly allocates the resources to one of those agents. Specifically, if agents $S\subseteq[n]$ bid for the resource, we allocate the resource to agent $i\in S$ with probability $p_{i}^{S}$. We call the $p_{i}^{S}$’s allocation probabilities and describe their values later (Theorem 4.2). Our formal mechanism can be found in Mechanism 1, where we also provide a little extra budget for each player, to ensure they do not run out with high probability when following our proposed equilibrium strategy, which we introduce next.
The strategy class we propose is a simple and natural strategy wherein each agent bids whenever their value is above a certain threshold. Adopting terminology from [fikioris2023online], we refer to this as a $\beta$-aggressive strategy.
Definition 4.1 ($\beta$-aggressive strategy).
Agent $i$ follows a $\beta$-aggressive strategy if she bids at time $t$ if and only if her value $V_{i}[t]$ is the top $\beta$-quantile of her value distribution [^1] If the top $\beta$-quantile is not well-defined, as when the value distribution has atoms, the agent can bid with probability $\rho(V_{i}[t])$ where $\rho:[0,\infty)\to[0,1]$ maximizes $\operatorname*{\mathbb{E}}_{V_{i}\sim\mathcal{F}i}[V_{i}\rho(V_{i})]$ subject to $\operatorname*{\mathbb{E}}_{V_{i}\sim\mathcal{F}_{i}}[\rho(V_{i})]\leq\beta$..
Intuitively, for an agent with fair share $\alpha$, following a $\alpha$-aggressive strategy is akin to truthfully reporting which are her $\alpha T$ favorite rounds. Building on this, we propose that each agent $i$ follows an $\alpha_{i}$-aggressive strategy. Note that with no competition, this would realize the agent’s fair share. Our first result is that everyone playing an $\alpha_{i}$-aggressive strategy is an approximate Nash equilibrium.
Proposition 4.1.
By setting $\delta_{i}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{i}T}}}$, no matter the choice of allocation probabilities $p_{i}^{S}$, each agent playing an $\alpha_{i}$-aggressive strategy is an $O(\sqrt{\nicefrac{{\log T}}{{T}}})$-approximate Nash equilibrium.
We defer the proof of this to Section 5. Note that the utility guarantees at this equilibrium depend on the choice of allocation probabilities $p_{i}^{S}$. Since the agents’ bids are i.i.d. in every round, we can calculate the fixed probability $p_{i}$ that agent $i$ will win the item conditioned on bidding (assuming everyone has budget remaining). This $p_{i}$ is called agent $i$’s interim allocation probability.
Definition 4.2 (Interim allocation probability).
Allocation probabilities $p_{i}^{S}$ induce interim allocation probabilities $p_{i}$ if $p_{i}$ is the probability that agent $i$ wins the item in a given round conditioned on agent $i$ bidding, and agents $j\neq i$ bidding independently with probability $\alpha_{j}$ each. Formally:
$p_{i}=\sum_{S\subseteq[n]:i\in S}p_{i}^{S}\left(\prod_{j\in S\setminus\{i\}}\alpha_{j}\right)\left(\prod_{j\in[n]\setminus S}(1-\alpha_{j})\right).$
Our goal is to maximize the fraction of ideal utility every agent is guaranteed in the above equilibrium. By definition of $p_{i}$, every agent in the equilibrium is guaranteed a $p_{i}$ fraction of their ideal utility. The following theorem shows that we can set $p_{i}^{S}$ to achieve the upper bound of Lemma 2.1; we present the proof (which is based on Border’s theorem) in Section 5.
Theorem 4.2.
There exist allocation probabilities $p_{i}^{S}$ such that the induced interim allocation probabilities $p_{i}$ are the same for all agents $i$. Specifically, we can make
$\forall i:\quad p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j}).$
Having achieved optimal equilibrium performance, we turn our attention to robustness. Unlike the equilibrium analysis, where we can use any allocation probabilities $p_{i}^{S}$ that induce the desired interim allocation probabilities $p_{i}$, under arbitrary competition, we must be more careful with our selection of $p_{i}^{S}$. In Section 6.2, we show that not carefully picking the $p_{i}^{S}$’s can lead to very poor robustness. Our main technical contribution is in Section 6.1, where we show that under careful upper bounds on the $p_{i}^{S}$’s we can guarantee strong robust performance without changing the $p_{i}$’s and thus the equilibrium performance. The following theorem summarizes our overall guarantees, subsuming the results we presented in this section, and the ones we prove in the next.
Theorem 4.3.
Consider the BRB Mechanism (Mechanism 1) with $\delta_{i}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{i}T}}}$. With a careful choice of allocation probabilities $p_{i}^{S}$ satisfying Theorem 4.2, we have the following:
  1. 1.
    Each player $i$ playing an $\alpha_{i}$-aggressive strategy is an $O(\sqrt{\nicefrac{{\log T}}{{T}}})$ Bayes-Nash equilibrium where, with probability $1-O(1/T^{2})$, player $i$ realizes utility
    $\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq\left(1-\prod_{j=1}^{n}(1-\alpha_{j})\right)v_{i}^{\star}-O\left(\sqrt{\frac{\log T}{T}}\right).$
  2. 2.
    Regardless of behavior of others, playing an $\alpha_{i}$-aggressive strategy, with probability $1-O(1/T^{2})$, gives player $i$ utility
    $\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq\left(\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2}\right)v_{i}^{\star}-O\left(\sqrt{\frac{\log T}{T}}\right).$
Note that this implies that each player gets a $\lambda_{\mathtt{NASH}}=(1-\prod_{j=1}^{n}(1-\alpha_{j}))-O(\sqrt{\nicefrac{{\log T}}{{T}}})$ fraction of their ideal utility in expectation at the approximate equilibrium, and the $\alpha_{i}$-aggressive strategy is $\lambda_{\mathtt{ROB}}=(\frac{1}{2}-O(\sqrt{\nicefrac{{\log T}}{{T}}}))$-robust. The robustness factor of $1/2$ matches the best current robustness factors [gorokh2019remarkable, fikioris2023online]  [^2] Through personal communication, we have learned of ongoing work proposing a different mechanism that achieves a higher robustness factor. However, this requires a much larger strategy space, and the robust strategy is more complicated.. We prove the equilibrium claim in Section 5 and the robustness claim in Section 6.
By our choice of $\delta_{i}^{T}$ in Theorem 4.3, the probability of any agent running out of budget is $O(1/T^{2})$ (similar to the probability of our bound not holding). Due to this, we could modify BRB’s budget constraint to allow agent $i$ to bid at most $\alpha_{i}(1+\delta_{i}^{t})t$ times by round $t\in[T]$ to get any time guarantees. Specifically, our utility bounds would hold for every round $t\in[T]$ with probability $1-O(1/\sqrt{t})$. Enforcing the budget constraint at each time also allows us to obtain an exact Nash equilibrium in the infinite time horizon. See Appendix B for details.
5 Equilibrium and Good Allocation Probabilities
5.1 Equilibrium
In this section, we prove our claim of Section 4 that each agent $i$ following the $\alpha_{i}$-aggressive strategy in the BRB Mechanism forms an approximate Nash Equilibrium. The proof sketch is very similar to the one for two agents in Section 3. As in Section 3, we consider a simpler game where budgets are enforced in expectation. While bounds in expectation are not enforceable, we focus on this case for simplicity of presentation, and only require that $\sum_{t=1}^{T}\mathbb{E}[b_{i}^{t}]\leq\alpha_{i}T$. Roughly speaking, we prove that in this game, the best response of agent $i$ when agents $j\neq i$ are following an $\alpha_{j}$-aggressive strategy is to follow an $\alpha_{i}$-aggressive strategy also. This, in turn, proves that these strategies form an equilibrium.
Lemma 5.1.
Fix an agent $i$ and assume all other agents $j\neq i$ are bidding i.i.d. $\mathrm{Bernoulli}(\alpha_{j})$. Suppose agent $i$ is trying to maximize her expected utility subject to the constraint that she does spend more than $\alpha_{i}T$ tokens in expectation; that is, she is choosing $(b_{i}^{t})_{t}$ to solve
$\max\,\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]\quad\text{subject to}\quad\sum_{t=1}^{T}\mathbb{E}[b_{i}^{t}]\leq\alpha_{i}T$
Her optimal strategy is to choose $(b_{i}^{t})_{t}$ that corresponds to bidding whenever her value is in the top $\alpha_{i}$-quantile of her value distribution, which yields agent $i$ expected utility
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]=p_{i}v_{i}^{\star}.$
The idea behind why the $\alpha_{i}$-aggressive strategy is the best response is as follows. At each round, the probability that agent $i$ wins conditioned on bidding is exactly their interim allocation probability $p_{i}$. The best solution for the agent is to bid in the $\alpha_{i}T$ rounds where the agent has the highest value. Therefore, agent $i$ should follow an $\alpha_{i}$-aggressive strategy, bidding whenever her value is in the top $\alpha_{i}$-quantile of her value distribution. We include the full proof of the lemma in Appendix D, where we also translate the result from this simplified game with budget constraints met in expectation to the actual game where budgets are a bit larger and strictly enforced.
5.2 Inducing Optimal Interim Allocation Probabilities
In this section, we show how to set the allocation probabilities $p_{i}^{S}$ to achieve optimal equilibrium performance in our BRB Mechanism. Specifically, in Theorem 4.2 we claimed that we can set each $p_{i}^{S}$ so that all interim allocation probabilities are equal: $p_{i}=1-\prod_{j}(1-\alpha_{j})$ for all $i$. This result follows from a special case of Border’s theorem, a theorem about the feasibility of reduced-form auctions. We present a special case of that theorem here and defer its full statement from the classical view of monetary auctions to Appendix C.
Theorem 5.2.
Given arbitrary numbers $p_{i}\in[0,1]$, there exist allocation probabilities $p_{i}^{S}$ such that the $p_{i}$ are interim allocation probabilities induced by the $p_{i}^{S}$’s if and only if
$\sum_{i\in[n]}p_{i}\alpha_{i}=1-\prod_{i\in[n]}(1-\alpha_{i})$ (2)
and for every $I\subseteq[n]$,
$\sum_{i\in I}p_{i}\alpha_{i}\leq 1-\prod_{i\in I}(1-\alpha_{i}).$ (3)
[…]
Figure 1: Flow network that can be used to prove Theorem 5.2. We let $S^{\prime}$ be the random set of bidding agents where each agent $i$ lies in $S^{\prime}$ independently with probability $\alpha_{i}$. Then, there are three kind of edges: edges whose flow corresponds to the probability of observing a specific $S^{\prime}$ (left), edges whose flow corresponds to how we randomly allocate the item condition on observing a specific $S^{\prime}$ (middle), and edges whose flow represent the probability that a specific agent gets the item (right). There is a flow of value $\Pr(S^{\prime}\neq\emptyset)$ if and only if there exist allocation probabilities $p_{i}^{S}$ inducing the interim allocation probabilities $p_{i}$. In other words, the flows $p_{i}^{S}\Pr(S^{\prime}=S)$ in the middle transform the probabilities that agents in a certain set $S$ bid to an agent $i\in S$ being allocated. We obtain the conditions in Theorem 5.2 by analyzing every minimum-cut of this network.
Now we can use Theorem 5.2 to prove Theorem 4.2.
Proof of Theorem 4.2.
We prove that the choice of $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ satisfies the conditions of Theorem 5.2. Observe that (2) holds with this choice of $p_{i}$. Also, observe that (3) holds for $I=\emptyset$. To show that (3) holds for $I\neq\emptyset$, define $a_{I}$ as
$a_{I}=\frac{1-\prod_{i\in I}(1-\alpha_{i})}{\sum_{i\in I}\alpha_{i}}.$
With this choice of $p_{i}$, (3) says
$\sum_{i\in I}p_{i}\alpha_{i}=\left(1-\prod_{j=1}^{n}(1-\alpha_{j})\right)\sum_{i\in I}\alpha_{i}\leq 1-\prod_{i\in I}(1-\alpha_{i}),$
which holds if and only if $a_{I}\geq 1-\prod_{j=1}^{n}(1-\alpha_{j})$. We shall show that $a_{I}$ is nonincreasing in $I$, which would suffice because $a_{[n]}=1-\prod_{j=1}^{n}(1-\alpha_{j})$. Take any $I\subseteq[n]$ containing some $k\in[n]$ where $I\setminus\{k\}\neq\emptyset$. Compute
$\begin{split}a_{I}-a_{I\setminus\{k\}}&=\frac{1-\prod_{i\in I}(1-\alpha_{i})}{\sum_{i\in I}\alpha_{i}}-\frac{1-\prod_{i\in I\setminus\{k\}}(1-\alpha_{i})}{\sum_{i\in I\setminus\{k\}}\alpha_{i}}\\ &=\frac{\left(1-\prod_{i\in I}(1-\alpha_{i})\right)\left(\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)-\left(1-\prod_{i\in I\setminus\{k\}}(1-\alpha_{i})\right)\left(\sum_{i\in I}\alpha_{i}\right)}{\left(\sum_{i\in I}\alpha_{i}\right)\left(\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)}\\ &=\frac{-\alpha_{k}+\prod_{i\in I\setminus\{k\}}(1-\alpha_{i})\left(\sum_{i\in I}\alpha_{i}-(1-\alpha_{k})\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)}{\left(\sum_{i\in I}\alpha_{i}\right)\left(\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)}\\ &=\frac{-\alpha_{k}+\prod_{i\in I\setminus\{k\}}(1-\alpha_{i})\alpha_{k}\left(1+\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)}{\left(\sum_{i\in I}\alpha_{i}\right)\left(\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)}\\ &=\frac{\alpha_{k}\left(-1+\prod_{i\in I\setminus\{k\}}(1-\alpha_{i})\left(1+\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)\right)}{\left(\sum_{i\in I}\alpha_{i}\right)\left(\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)}.\end{split}$
This can be seen to be nonpositive by the identities that $1-x\leq e^{-x}$, so $\prod_{i\in I\setminus\{k\}}(1-\alpha_{i})\leq\exp\left(-\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)$, and also $e^{-y}(1+y)\leq 1$ applied to $y=\sum_{i\in I\setminus\{k\}}\alpha_{i}$. ∎
6 Robustness
In Section 5, we prove that each player playing an $\alpha_{i}$-aggressive strategy is an approximate equilibrium. To complete the proof of Theorem 4.3, we show how to choose the allocation probabilities $p_{i}^{S}$ to guarantee robustness while maintaining the same interim allocations. To do this, we strengthen Border’s Theorem to also satisfy some additional properties that guarantee robustness. We then show that this strengthening is necessary, as many allocations obtained by the standard Border’s Theorem are not robust. Finally, we show our robustness result is tight in that no matter the choice of $(p_{i}^{S})$, it is not possible to guarantee each agent a $\lambda$-robust strategy for $\lambda$ much greater than $1/2$.
6.1 Achieving a $1/2$ Robustness Factor
We will show how to choose allocation probabilities $p_{i}^{S}$ to make an $\alpha_{i}$-aggressive strategy $(\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2})$-robust while still inducing the interim allocation probabilities $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ as in Theorem 4.2, proving Theorem 4.3.
We now consider (collusive) strategies that agents $j\neq i$ can employ to minimize agent $i$’s utility. Say that agent $i$ is blocked when she bids but does not receive the item. We will focus on how many bids the other agents have to make each time they block agent $i$. Let us first consider the case where other agents never bid two at a time. Conditioned on only agent $j$ bidding, agent $i$ gets blocked with probability $\alpha_{i}p_{j}^{\{i,j\}}$, with $\alpha_{i}$ being the probability that agent $i$ bids and $p_{j}^{\{i,j\}}$ being the probability that agent $j$ wins when they both bid. We have to ensure that this probability is not too large. Specifically, assume that for some $\bar{p}$ it holds $p_{j}^{\{i,j\}}\leq\bar{p}$ for all $j\neq i$. Thus, when one other agent bids at a time, the other agents must spend $(\bar{p}\alpha_{i})^{-1}$ tokens each time they block agent $i$ in expectation. When 2 or more agents bid, we still have that agent $i$ bids only with probability $\alpha_{i}$, so the expected number of tokens spent for blocking agent $i$ is at least $2\alpha_{i}^{-1}$. As long as $\bar{p}\geq 1/2$, this is less effective from a bang-per-buck perspective.
Since the number of times agents $j\neq i$ can bid is at most $\sum_{j\neq i}\alpha_{j}=(1-\alpha_{i})T$, the expected number of times agent $i$ can get blocked is at most $\max(\bar{p},1/2)\alpha_{i}(1-\alpha_{i})T$. Combining this with the fact that agent $i$ bids $\alpha_{i}T$ times in expectation, we get that she will win the item at least
$\alpha_{i}T-\max(\bar{p},1/2)\alpha_{i}(1-\alpha_{i})T=\quantity\big(1-\max(\bar{p},1/2)(1-\alpha_{i}))\alpha_{i}T$
times in expectation. Therefore, to ensure that agent $i$ does not get too low utility, we have to ensure that $\bar{p}$ is not too large.
The formal statement of the above argument is the lemma below. Its proof, which we defer to Appendix D, uses martingale concentration arguments to ensure that agent $i$ obtains at least $(1-(1-\alpha_{i})\bar{p})$ fraction of her ideal utility with high probability.
Lemma 6.1.
Fix an agent $i$. Given allocation probabilities $(p_{k}^{S})$, if $p_{j}^{\{i,j\}}\leq\bar{p}$ for every other agent $j$ where $\bar{p}\geq 1/2$, then when we run BRB with slack parameters $\delta_{i}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{i}T}}}$, an $\alpha_{i}$-aggressive strategy guarantees agent $i$ utility
$\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq\left(1-\bar{p}(1-\alpha_{i})\right)-O\left(\sqrt{\frac{\log T}{T}}\right)$
with probability at least $1-O(1/T^{2})$ regardless of the behavior of agents $j\neq i$.
We want to give allocation probabilities $p_{i}^{S}$ that induce interim allocation probabilities $p_{i}=1-\prod_{k}(1-\alpha_{k})$ that satisfy the above lemma with a low value of $\bar{p}$ to obtain simultaneous equilibrium and robustness guarantees as in Theorem 4.3. Let us now demonstrate that this is possible if each agent has equal fair share $\alpha_{i}=1/n$. Choose the symmetric probabilities $p_{i}^{S}=1/|S|$ for each $i\in S$, i.e., give the item to a bidding agent uniformly at random. By symmetry, these do indeed induce the interim allocation probabilities $p_{i}=1-\prod_{k}(1-\alpha_{k})=1-(1-1/n)^{n}$. They satisfy $p_{j}^{\{i,j\}}=1/2$, so we can use Lemma 6.1 with $\bar{p}=1/2$, yielding a robustness factor of $1-\bar{p}(1-\alpha_{i})=1/2+1/(2n)$.
It would be natural to try to generalize the use of $\bar{p}=1/2$ to the assymetric case. However, this would be too low, and the desired interim allocation of Theorem 4.3 may not be feasible with this restriction. To see why, it is useful to recall the two agent special case, when the agent with fair share $1-\alpha$ had to be allocated with probability $\alpha$, which can be more than $\bar{p}=1/2$. We’ll end up using this lemma with $\bar{p}=\frac{1+\alpha_{i}}{2}$ (always satisfying $\alpha_{i}\leq\bar{p}$), which proves the robustness guarantee of Theorem 4.3. The remaining technically challenging part of the proof is to show that there is a set of allocation probabilities $p_{i}^{S}$ satisfying these conditions.
The outline of the rest of the section is as follows. To prove that the desired allocation probabilities exist, consider the network of Fig. 1 used to prove Border’s theorem. Border’s theorem does not cover the upper bound constraint on $p_{j}^{\{i,j\}}$ of Lemma 6.1. To ensure this limit, we will change the capacity of the edges from node $u_{\{i,j\}}$ to node $v_{j}$ for all agents $j$ and all two-element sets $\{i,j\}$ to $\frac{1+\alpha_{i}}{2}$ as illustrated in Fig. 2.
In Lemma 6.2 we establish the condition for a flow of value $1-\prod_{k}(1-\alpha_{k})$ to exist with additional capacities on the edges that have infinite capacity in the original construction. While the condition is more complex, its proof is analogous to the proof of Border’s theorem discussed above: we need to make sure that each cut in the network has capacity high enough to support the desired flow amount. The most technically challenging and surprising part of our proof is establishing that the established condition actually holds in Lemma 6.3. This directly implies Theorem 4.3 completing the proof of the existence of allocation probabilities that simultaneously have optimal Nash equilibrium and the desired robustness.
While the below condition uses our upper bounds $p_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$ specifically, we can prove a more general condition with arbitrary upper bounds on any $p_{i}^{S}$, $S\subseteq[n],i\in S$. Because that condition is more complex, we present and prove it in Appendix C.
Lemma 6.2.
There exists $p_{i}^{S}$’s that induce interim allocation probabilities $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ and satisfy the upper bounds $p_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$ if and only if
$\left(1-\prod_{k=1}^{n}(1-\alpha_{k})\right)\sum_{i\in I}\alpha_{i}+\prod_{i\in I}(1-\alpha_{i})+\frac{1}{2}\prod_{k=1}^{n}(1-\alpha_{k})\sum_{i\in I}\frac{\alpha_{i}}{1-\alpha_{i}}\sum_{j\notin I}\alpha_{j}\leq 1$ (4)
for every $I\subseteq[n]$.
Proof.
Create an $s$-$t$ flow network as follows. Let $\mu$ be the probability distribution over subsets $S^{\prime}\subseteq[n]$ where each $i\in S^{\prime}$ independently with probability $\alpha_{i}$. For each nonempty $S\subseteq[n]$ create a node $u_{S}$ and connect it with an edge $(s,u_{S})$ to the source node with capacity
$c(s,u_{S})=\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S).$
For each $i$, create a node $v_{i}$. For each $S\neq\emptyset$ such that $i\in S$, add an edge $(u_{S},v_{i})$ with capacity
$c(u_{S},v_{i})=\begin{cases}\infty&\text{if $|S|\neq 2$}\\ \frac{1+\alpha_{j}}{2}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)&\text{if $S=\{i,j\}$}\end{cases}.$
Also, add an edge $(v_{i},t)$ with capacity
$c(v_{i},t)=p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime}).$
The flow network is depicted in Fig. 2.
[…]
Figure 2: Flow network used in the proof of Lemma 6.2. The flow network is very similar to the one in Fig. 1 used to prove Theorem 5.2. However, the red edges $(u_{\{i,j\}},v_{j})$ have explicit capacities as opposed to infinite capacity to enforce the additional bounds $\bar{p}_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$.
Note that with $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$,
$\sum_{i\in[n]}p_{i}\alpha_{i}=1-\prod_{i\in[n]}(1-\alpha_{i}).$ (5)
The cut with $s$ on one side and everything else on the other has capacity
$\sum_{S\subseteq[n]:S\neq\emptyset}c(s,u_{S})=\sum_{S\subseteq[n]:S\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\neq\emptyset)=1-\prod_{i\in[n]}(1-\alpha_{i}).$
The cut with $t$ on one side and everything else on the other has capacity
$\sum_{i\in[n]}c(v_{i},t)=\sum_{i\in[n]}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})=\sum_{i\in[n]}p_{i}\alpha_{i}=1-\prod_{i\in[n]}(1-\alpha_{i}),$
using (5) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ satisfying upper bounds $p_{i}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\alpha_{i})$, it suffices to show that (4) holds only if there is a feasible flow of flow value equal to this cut capacity.
Take any minimum-capacity $s$-$t$ cut $(A,B)$. Let $I=\{i\in[n]:v_{i}\in B\}$. We now argue that we can determine which side of the minimum-capacity cut $(A,B)$ the rest of the nodes are on just based on $I$.
  • If $i\in I$, then we must have $u_{S}\in B$ for $|S|\neq 2$ since those edges $(u_{S},v_{i})$ have infinite capacity.
  • For any set $S$, if $i\notin I$ for every $i\in S$, then we can assume $u_{S}\in A$ since there are no edges coming out of $u_{S}$ except the $(u_{S},v_{i})$.
  • For a doubleton set $\{i,j\}$, if $i\in I$ and $j\notin I$, then the edge capacity $c(s,u_{\{i,j\}})=\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$ coming in is larger than the edge capacity $c(u_{\{i,j\}},v_{i})=\frac{1+\alpha_{j}}{2}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$ going out, so $u_{\{i,j\}}\in A$.
  • For a doubleton set $\{i,j\}$, if both $i,j\in I$, then the edge capacity $c(s,u_{\{i,j\}})=\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$ coming in is smaller than the sum of the edge capacities
    $c(u_{\{i,j\}},v_{i})+c(u_{\{i,j\}},v_{j})=\left(\frac{1+\alpha_{i}}{2}+\frac{1+\alpha_{j}}{2}\right)\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$
    going out, so $u_{\{i,j\}}\in B$.
Now we can compute the total capacity of the cut $(A,B)$:
$\begin{split}c(A,&B)=\sum_{S:|S|\neq 2,S\cap I\neq\emptyset}c(s,u_{S})+\sum_{i,j\in I:i\neq j}c(s,u_{\{i,j\}})+\sum_{i\in I}\sum_{j\notin I}c(u_{\{i,j\}},v_{i})+\sum_{i\notin I}c(v_{i},t)\\ &=\sum_{S:S\cap I\neq\emptyset}c(s,u_{S})-\sum_{i\in I}\sum_{j\notin I}c(s,u_{\{i,j\}})+\sum_{i\in I}\sum_{j\notin I}c(u_{\{i,j\}},v_{i})+\sum_{i\notin I}c(v_{i},t)\\ &=\sum_{S:S\cap I\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)-\sum_{i\in I}\sum_{j\notin I}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=\{i,j\})\left(1-\frac{1+\alpha_{j}}{2}\right)+\sum_{i\notin I}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})\\ &=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\cap I\neq\emptyset)-\sum_{i\in I}\sum_{j\notin I}\alpha_{i}\alpha_{j}\prod_{k\notin\{i,j\}}(1-\alpha_{k})\left(1-\frac{1+\alpha_{j}}{2}\right)+\sum_{i\notin I}p_{i}\alpha_{i}\\ &=1-\prod_{i\in I}(1-\alpha_{i})-\sum_{i\in I}\sum_{j\notin I}\alpha_{i}\alpha_{j}\prod_{k\notin\{i,j\}}(1-\alpha_{k})\left(1-\frac{1+\alpha_{j}}{2}\right)+1-\prod_{i\in I}(1-\alpha_{i})-\sum_{i\in I}p_{i}\alpha_{i}\\ &=1-\prod_{i\in I}(1-\alpha_{i})-\frac{1}{2}\prod_{k=1}^{n}(1-\alpha_{k})\sum_{i\in I}\frac{\alpha_{i}}{1-\alpha_{i}}\sum_{j\notin I}\alpha_{j}\\ &\quad+1-\prod_{i\in I}(1-\alpha_{i})-\left(1-\prod_{k=1}^{n}(1-\alpha_{k})\right)\sum_{i\in I}\alpha_{i}\end{split}$
using (5) for the second-to-last equality and substituting $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ and doing some algebraic rearrangement for the last equality. Rearranging, (4) is equivalent to the above being least $1-\prod_{i\in[n]}(1-\alpha_{i})$. The result follows from the max-flow min-cut theorem. ∎
Next, we prove that allocation probabilities satisfying Lemma 6.2 do exist. The key observation to showing that (4) holds is noticing that the left-hand side of the inequality is Schur-concave in the variables $\alpha_{i}$ for $i\in I$ and Schur-convex in the variables $\alpha_{j}$ for $j\not\in I$. Using properties of Schur-convex and Schur-concave functions, it suffices to check the inequality for the special case when $\alpha_{i}$ is the same for all $i\in I$ and there is only one non-zero $\alpha_{j}$ outside of the set $I$. Fig. 3 plots the maximum possible value of the left hand side of (4), depending on the $\alpha$ values. We emphasize how close it gets to the bound of $1$ with sets $I$ of size $5$ or higher.
Lemma 6.3.
For any vector of agents’ fair shares $(\alpha_{i})$, for every $I\subseteq[n]$,
$\left(1-\prod_{k=1}^{n}(1-\alpha_{k})\right)\sum_{i\in I}\alpha_{i}+\prod_{i\in I}(1-\alpha_{i})+\frac{1}{2}\prod_{k=1}^{n}(1-\alpha_{k})\sum_{i\in I}\frac{\alpha_{i}}{1-\alpha_{i}}\sum_{j\notin I}\alpha_{j}\leq 1.$ (6)
Proof.
Notice that if $I=\emptyset$ or $I=[n]$, then this is indeed true, so assume $\emptyset\subsetneq I\subsetneq[n]$. Let $X=\sum_{i\in I}\alpha_{i}$ and
$K=\left\{((x_{i})_{i\in I},(y_{j})_{j\notin I})\in[0,1]^{I}\times[0,1]^{[n]\setminus I}:\sum_{i\in I}x_{i}=X,\sum_{j\notin I}y_{j}=1-X\right\}.$
Define the function $f:K\to\mathbb{R}$ by
$\begin{split}f((x_{i})_{i\in I},&(y_{j})_{j\notin I})=\left(1-\prod_{i\in I}(1-x_{i})\prod_{j\notin I}(1-y_{j})\right)X\\ &+\prod_{i\in I}(1-x_{i})+\frac{1}{2}\left(\prod_{i\in I}(1-x_{i})\right)\left(\prod_{j\notin I}(1-y_{j})\right)\left(\sum_{i\in I}\frac{x_{i}}{1-x_{i}}\right)(1-X).\end{split}$
The left-hand side of (6) is precisely $f((\alpha_{i})_{i\in I},(\alpha_{j})_{j\notin I})$, so it suffices to show that the maximum of $f$ is at most $1$. By taking derivatives, we check in Appendix D that $f(\cdot,(y_{j})_{j\notin I})$ is Schur-concave for each $(y_{j})_{j\notin I}$ and that $f((x_{i})_{i\in I},\cdot)$ is Schur-convex for each $(x_{i})_{i\in I}$. Therefore, the maximum of $f$ occurs at $((x_{i}^{\star})_{i\in I},(y_{j}^{\star})_{j\notin I})$ when each $x_{i}^{\star}=X/m$ where $m=|S|$ and there is a single $y_{j^{\star}}^{\star}=1-X$ and all other $y_{j}^{\star}=0$. In that case,
$f((x_{i}^{\star})_{i\in I},(y_{j}^{\star})_{j\notin I})$ $\leq\left(1-\frac{X^{2}(mX+m-2X)}{2(m-X)}\right)\left(1-\frac{X}{m}\right)^{m}+X=1-(1-X)(1-g(X))$
where
$g(X)=\frac{\left(1-\frac{X}{m}\right)^{m}\left(mX^{2}+2mX+2m-2X^{2}-2X\right)}{2(m-X)}.$
We plot $f$ in Fig. 3. It is easy to verify that $g(0)=1$, and hence if we set $X=0$ or $X=1$, then the above bound is $1$. To complete the proof, we claim that $g^{\prime}(X)\leq 0$ (i.e., $g(X)$ is decreasing) in $[0,1]$. This is indeed the case, as one can readily check
$g^{\prime}(X)=-\frac{X\left(1-\frac{X}{m}\right)^{m}\left(mX(m-1)+2(m-X)\right)}{2(m-X)^{2}}\leq 0.$
Figure 3: Plot of $f(X)$, the maximum value of the left-hand side of (6) for a fixed size $m$ of $I$. It gets super close to $1$, and we prove that it never exceeds $1$.
Lemmas 6.2 and 6.3 establish the existence of allocation probabilities $p_{i}^{S}$ inducing interim allocation probabilities $p_{i}=1-\prod_{k=1}^{n}(1-\alpha_{k})$ that also satisfy upper bounds $p_{j}^{\{i,j\}}\leq\frac{1+\alpha_{i}}{2}$. The interim allocation probabilities show the utility guarantee at equilibrium of Theorem 4.3. By Lemma 6.1 with $\bar{p}=\frac{1+\alpha_{i}}{2}$, the upper bounds prove the robustness claim of Theorem 4.3.
6.2 Interim Allocation Probabilities Do Not Guarantee Robustness
In this section, we examine if the upper-bound restrictions on $p_{j}^{\{i,j\}}$ of Lemma 6.1 are necessary for the robustness guarantee or just specifying the interim allocations is enough. We will show that these restrictions are indeed necessary: we construct an example where the allocation probabilities $(p_{i}^{S})$ induce the desired interim probabilities of Theorem 4.2 but some agent $i$ cannot guarantee more than an $2\alpha_{i}$ fraction of her ideal utility robustly.
Consider the symmetric case where each $\alpha_{i}=1/n$. For each $S$ such that $i\in S$, let
$p_{i}^{S}=\begin{cases}1/|S|&\text{if $1\notin S$}\\ 1&\text{if $|S|\neq 2$ and $i=1$}\\ 0&\text{if $|S|\neq 2$ and $1\in S$ and $i\neq 1$}\\ 1/n&\text{if $|S|=2$ and $i=1$}\\ 1-1/n&\text{if $|S|=2$ and $i\neq 1$}\end{cases}.$ (7)
In words, we set the allocation probabilities such that if agent $1$ does not bid, we allocate uniformly at random, and if agent $1$ bids and more than one other agent bids, we allocate to agent $1$ with probability $1$, and if agent $1$ bids and exactly one other agent bids, we allocate to agent $1$ with probability $1/n$. We carefully picked these allocation probabilities such that they induce the same interim allocation probabilities $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ as in Theorem 4.2. In addition, the probabilities violate the restriction of Lemma 6.1 as much as possible: $p_{j}^{\{1,j\}}=1-1/n\gg 1/2$ when $n$ is large.
Even though the $p_{i}^{S}$ induce the correct interim allocation probabilities, they do not guarantee robustness by the following proposition, proved formally in Appendix D.
Proposition 6.4.
Using the allocation probabilities $p_{i}^{S}$ as defined by (7), player $1$ does not have a $\lambda$-robust strategy for any $\lambda>\frac{n-1}{n^{2}}+\frac{1}{n}-O\left(\sqrt{\nicefrac{{\log T}}{{T}}}\right)$.
Proof sketch.
Suppose other players $i\neq 1$ take turns bidding, exactly one agent $i\neq 1$ bidding per round. Since they have total budget a $\frac{n-1}{n}$ fraction of the rounds, they can do this for all rounds but $T/n$. By how the $p_{i}^{S}$’s are set, agent $1$ can only win the item with probability $1/n$ on these rounds. This results in agent $1$’s utility to be extremely low. ∎
6.3 A $1/2$ Hardness Result
In this section, we show that our analysis for the robustness bound of Theorem 4.3 is approximately tight if the fair shares are small. Specifically, we show that the BRB mechanism cannot guarantee every agent more than a 1/$21/2+\sum_{i}\alpha_{i}^{2}/$ 2 fraction of her ideal utility. We show this for any allocation probabilities $p_{i}^{S}$ that the mechanism could use.
The following lemma stems from the strategy used by agents $j\neq i$ bidding in an anti-correlated fashion. They make sure that at most $1$ agent $j\neq i$ bids at each time by coordinating with each other. Each agent $j\neq i$ can bid for an $\alpha_{j}$ fraction of the time. Agents $j\neq i$ will not bid for a $1-\sum_{j\neq i}\alpha_{j}=\alpha_{i}$ fraction of the time, so agent $i$ can win these rounds with no competition. At each round that agent $j$ bids, agent $i$ can only win with probability $p_{i}^{\{i,j\}}$ conditioned on agent $i$ bidding. Summing the probabilities that agent $i$ can win conditioned on bidding over all times, we obtain the $\alpha_{i}+\sum_{j\neq i}\alpha_{j}p_{i}^{\{i,j\}}$ term below.
Lemma 6.5.
When the slack parameters are set as $\delta_{k}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\alpha_{k}T}}}$, no matter what value distributions agent $i$ has, if agent $i$has a $\lambda$-robust strategy has a $\lambda_{i}$-robust strategy, then
$\lambda_{i}\leq\alpha_{i}+\sum_{j\neq i}\alpha_{j}p_{i}^{\{i,j\}}+O\left(\sqrt{\frac{\log T}{T}}\right).$
We formally prove the lemma in Appendix D. We can use the bound in the above lemma to show that there is always an agent that cannot have a $\lambda$-robust strategy for $\lambda$ not much higher than $1/2$.
Theorem 6.6.
When the slack parameters are set as $\delta_{k}^{T}=\sqrt{\nicefrac{{6\ln T}}{{\beta_{k}T}}}$, no matter what value distributions the agents have, if every agent $i$ has a $\lambda$-robust strategy, then
$\lambda\leq\frac{1}{2}+\frac{1}{2}\sum_{i=1}^{n}\alpha_{i}^{2}+O\left(\sqrt{\frac{\log T}{T}}\right).$
The theorem can be proved using Lemma 6.5 by summing the terms of that lemma. The full proof is included in Appendix D.
[…]
Appendix A Better Equilibrium Guarantees
We designed the Budgeted Robust Border mechanism to guarantee each agent a $1-\prod_{j=1}^{n}(1-\alpha_{j})$ fraction of their $\alpha_{i}$-ideal utility. While this may be the best factor with worst-case value distributions, there are allocation procedures that can do a lot better with other value distributions [chido, fikioris2023online].
We present a very simple way to generalize our mechanism to achieve better equilibrium guarantees when the value distributions are not worst case. We generalize this mechanism by parameterizing it in two ways.
  1. 1.
    Set the per-round budgets to a parameter $\beta_{i}$. That is, instead of giving everyone $\alpha_{i}(1+\delta_{i}^{T})T$ tokens, we give them $\beta_{i}(1+\delta_{i}^{T})T$ tokens where $\beta_{i}$ is a parameter.
  2. 2.
    Instead of always choosing the allocation probabilities $p_{i}^{S}$ in a very specific way as in Theorem 4.2, let the $p_{i}^{S}$’s be parameters.
We formally give this parameterized mechanism in Mechanism 2.
Input: Per-round budgets $\beta_{i}\in[0,1]$ and slack parameters $\delta_{i}^{T}$
Input: Allocation probabilities $(p_{i}^{S})_{\begin{subarray}{c}i\in[n]\\ S\subseteq[n]\end{subarray}}$ where $(p_{i}^{S})_{i\in[n]}$ is a probability distribution over agents $i\in S$
Endow each agent with $B_{i}[1]=\beta_{i}(1+\delta_{i}^{T})T$ bid tokens;
for $t=1,2,\dots,T$ do
      Agents submit bids $b_{i}^{t}\in\{0,1\}$;
      Budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $B_{i}[t]\leq 0$;
      Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents;
      A winner $i^{t}$ is randomly selected from $S[t]$ according to $(p_{i}^{S[t]})_{i\in[n]}$;
      Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$;
      end for
MECHANISM 2 Generalized Budgeted Border
Regardless of choice of $\beta_{i}$’s and $p_{i}^{S}$’s, there is a similar equilibrium as in Proposition 4.1 and similar equilibrium utility guarantee as in Theorem 4.3. Before stating this formally, let us define some useful terminology. First, we generalize the notion of ideal utility. Adopting terminology from [fikioris2023online], the $\beta$-ideal utility of an agent $i$ is the maximum expected utility they can obtain from a single round if they can obtain the item simply by requesting it, but they are only allowed to request it with probability at most $\beta$. Formally, the (per-round) $\beta$-ideal utility is the following.
Definition A.1 ($\beta$-ideal utility, $\beta$-ideal utility probability function).
Agent $i$’s $\beta$-ideal utility is the value of the following maximization problem over measurable $\rho_{i}:[0,\infty)\to[0,1]$:
$\begin{split}v_{i}^{\star}(\beta)=\max\,\operatorname*{\mathbb{E}}_{V_{i}\sim\mathcal{F}_{i}}[V_{i}\rho_{i}(V_{i})]\quad\text{subject to}\quad\operatorname*{\mathbb{E}}_{V_{i}\sim\mathcal{F}_{i}}[\rho_{i}(V_{i})]\leq\beta.\end{split}$
We call the optimal solution $\rho_{i}$ to the above optimization problem agent $i$’s $\beta$-ideal utility probability function, which we denote by $(\rho_{i}^{\beta})^{\star}$.
Note that the ideal utility, as we previously defined it in Definition 2.1, is just the $\alpha_{i}$-ideal utility.
Recall the notion of $\beta$-aggressive strategy from Definition 4.1, where an agent bids when her value is in the top $\beta$-quantile of her value distribution (subject to her budget constraint). We can also rephrase a $\beta$-aggressive strategy as bidding with probability $(\rho_{i}^{\beta})^{\star}(V_{i}[t])$ (subject to the budget constraint).
For the same reason as in BRB, each agent $i$ playing a $\beta_{i}$-aggressive strategy is an approximate Nash equilibrium. We formally prove the following in Appendix D.
Proposition A.1.
By setting $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$, no matter the choice of allocation probabilities $p_{i}^{S}$, each player $i$ playing a $\beta_{i}$-aggressive strategy is an $O\left(\sqrt{\frac{\log T}{T}}\right)$-approximate Nash equilibrium.
To discuss the utility guarantee at this approximate equilibrium, we define interim allocation probabilities similar to Definition 4.2, except that the interim allocation probabilities $p_{i}$ depend on the per-round budgets $\beta_{i}$. The previous definition of interim allocation probability is a special case where $\beta_{i}=\alpha_{i}$.
Definition A.2.
Fix per-round bid rates $\beta_{i}$. Allocation probabilities $p_{i}^{S}$ induce interim allocation probabilities $p_{i}$ if $p_{i}$ is the probability that agent $i$ wins the item in a given round conditioned on agent $i$ bidding that round and agents $j\neq i$ bidding independently with probability $\beta_{j}$ each. Formally, this means that
$p_{i}=\sum_{S\subseteq[n]:i\in S}p_{i}^{S}\left(\prod_{j\in S\setminus\{i\}}\beta_{j}\right)\left(\prod_{j\in[n]\setminus S}(1-\beta_{j})\right).$
We now obtain a similar equilibrium utility guarantee as in Theorem 4.3 that we prove formally in Appendix D.
Theorem A.2.
By setting $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$, at the approximate equilibrium where each player $i$ plays a $\beta_{i}$-aggressive strategy, with probability at least $1-O(1/T^{2})$, player $i$ gets utility
$\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq p_{i}v_{i}^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log T}{T}}\right).$
As before, Border’s Theorem gives a clean characterization of which numbers $p_{i}\in[0,1]$ can actually be induced by some allocation probabilities $p_{i}^{S}$. Specifically, the below theorem follows from Border’s Theorem and is analogous to Theorem 5.2, which we explain in Appendix C.
Theorem A.3.
Given per-round budgets $\beta_{i}$ and probabilities $p_{i}$, there exist allocation probabilities $p_{i}^{S}$ such that the $p_{i}$’s are interim allocation probabilities induced by the allocation probabilities $p_{i}^{S}$ if and only if
$\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$ (8)
and for any $I\subseteq[n]$,
$\sum_{i\in I}p_{i}\beta_{i}\leq 1-\prod_{i\in I}(1-\beta_{i}).$ (9)
In general, the principal choose $\beta_{i}$’s and $p_{i}$’s in any way they like that satisfies Theorem A.3. Let’s give a particular suggestion. Suppose we set the $\beta_{i}$’s to be proportional to the fair shares $\alpha_{i}$. We can obtain the following using Theorem A.3, which has a very similar proof to that of Theorem 4.2 that we formally give in Appendix D.
Corollary A.4.
Let $0<\gamma\leq\min_{i\in[n]}1/\alpha_{i}$ and let $\beta_{i}=\gamma\alpha_{i}$. There exist allocation probabilities $(p_{i}^{S})$ that induce the interim allocation probabilities
$p_{i}=\frac{1-\prod_{j\in[n]}(1-\gamma\alpha_{j})}{\gamma}.$
These $p_{i}$ satisfy
$p_{i}\geq\frac{1-e^{-\gamma}}{\gamma}.$
Using the above $p_{i}^{S}$’s in our mechanism, at the equilibrium, each agent $i$ gets expected per-round utility at least $\frac{1-e^{-\gamma}}{\gamma}v^{\star}(\gamma\alpha_{i})$. The function $\frac{1-e^{-\gamma}}{\gamma}$ is decreasing in $\gamma$, while $v^{\star}(\gamma\alpha_{i})$ is nondecreasing in $\gamma$, so if the principal wants to use the above corollary, they could set $\gamma$ appropriately based on the agents’ value distributions to maximize the $\frac{1-e^{-\gamma}}{\gamma}v^{\star}(\gamma\alpha_{i})$. Setting $\gamma=1$ recovers the old mechanism BRB. However, other choices of $\gamma$ may be better if the agents had specific value distributions. For example, the below corollary proved in Appendix D shows that in the symmetric agent case with $\mathrm{Uniform}([0,1])$ value distributions, we can set $\gamma$ such that each agent gets almost all of their $\alpha_{i}$-ideal utility, a result similar to the equilibrium guarantee in [chido].
Corollary A.5.
Suppose each agent has fair share $\alpha_{i}=\frac{1}{n}$ and a $\mathrm{Uniform}([0,1])$ value distribution. Then, by setting $\beta_{i}=\gamma\alpha_{i}$ where $\gamma=\Theta(\log n)$, there exist allocation probabilities $(p_{i}^{S})$ inducing interim allocation probabilities $p_{i}$ such that
$p_{i}v_{i}^{\star}(\beta_{i})\geq\left(1-O\left(\frac{\log n}{n}\right)\right)v_{i}^{\star}.$
Appendix B Any Time Guarantees and Exact Nash Equilibrium in an Infinite Time Horizon
B.1 Any Time Guarantees
We can obtain any time utility guarantees by enforcing the budget constraint at any time, i.e., agent $i$ can only bid at most $\alpha_{i}(1+\delta_{i}^{t})t$ times by round $t$ instead of just at end time $T$. We formally describe this mechanism in Mechanism 3 (based off of Generalized Budgeted Border in Appendix A with general per-round budgets $\beta_{i}$).
Input: Per-round budgets $\beta_{i}\in[0,1]$ and slack parameters $\delta_{i}^{t}$
Input: Allocation probabilities $(p_{i}^{S})_{\begin{subarray}{c}i\in[n]\\ S\subseteq[n]\end{subarray}}$ where $(p_{i}^{S})_{i\in[n]}$ is a probability distribution over agents $i\in S$
for $t=1,2,\dots,T$ do
      Agents submit bids $b_{i}^{t}\in\{0,1\}$;
      Budgets are enforced: $b_{i}^{t}\leftarrow 0$ for each $i$ such that $\sum_{s=1}^{t-1}b_{i}^{s}\geq\beta_{i}(1+\delta_{i}^{t})t$;
      Let $S[t]=\{i:b_{i}^{t}=1\}$ be the set of bidding agents;
      A winner $i^{t}$ is randomly selected from $S[t]$ according to $(p_{i}^{S[t]})_{i\in[n]}$;
      Budgets get updated: $B_{i}[t+1]=B_{i}[t]-b_{i}^{t}$ for every agent $i$;
      end for
MECHANISM 3 Any Time Budgeted Border
Formally, we obtain the following theorems, proved in Appendix D.
Theorem B.1.
By running Any Time Budgeted Border with $\delta_{i}^{t}=\sqrt{\frac{6\ln t}{\beta_{i}t}}$, each agent playing a $\beta_{i}$-aggressive strategy is an $O\left(\sqrt{\frac{\log T}{T}}\right)$-approximate equilibrium. At this equilibrium, at each time $t$, with probability at least $1-O(1/\sqrt{t})$, player $i$ gets utility
$\frac{1}{t}\sum_{s=1}^{t}U_{i}[s]\geq p_{i}v_{i}^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log t}{t}}\right).$
(Here, $p_{i}$ is agent $i$’s interim allocation probability with per-round budgets $(\beta_{k})$ as defined in Definition A.2.)
Theorem B.2.
By running Any Time Budgeted Border with $\beta_{i}=\alpha_{i}$, $\delta_{i}^{t}=\sqrt{\frac{6\ln t}{\alpha_{i}t}}$, and allocation probabilities as guaranteed by Lemmas 6.2 and 6.3, if agent $i$ plays an $\alpha_{i}$-aggressive strategy, regardless of the behavior of other agents $j\neq i$, with probability $1-O(1/\sqrt{t})$, at each time $t$, agent $i$ will obtain utility
$\frac{1}{t}\sum_{s=1}^{t}U_{i}[s]\geq\left(\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2}\right)v_{i}^{\star}-O\left(\sqrt{\frac{\log t}{t}}\right).$
B.2 Exact Nash Equilibrium in an Infinite Time Horizon
We show that by running Any Time Budgeted Border in an infinite time horizon, we can make the approximate Nash equilibrium in Theorem 4.3 into an exact Nash equilibrium. Each agent will maximize
$\mathbb{E}\left[\liminf_{t\to\infty}\frac{1}{t}\sum_{s=1}^{t}U_{i}[s]\right]=\liminf_{t\to\infty}\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s]],$
observing uniform integrability for the swap of limit and expectation.
We prove the following theorem in Appendix D.
Theorem B.3.
By running Any Time Budgeted Border with $\beta_{i}=\alpha_{i}$, $\delta_{i}^{t}=\sqrt{\frac{6\ln t}{\alpha_{i}t}}$, each player $i$ playing a $\beta_{i}$-aggressive strategy is a Nash equilibrium under which
$\frac{1}{t}\sum_{s=1}^{t}U_{i}[s]\to p_{i}v^{\star}(\beta_{i})$
almost surely as $t\to\infty$.
Appendix C Border’s Theorem
C.1 Obtaining Theorems 5.2 and A.3 via a Reduction from Border’s Theorem
We claimed that Theorem 5.2 and more generally, Theorem A.3, is a special case of Border’s Theorem. We shall detail Border’s Theorem and argue why our problem is a special case. Border’s Theorem deals with the setting of selling a single item to bidders via a direct-revelation mechanism. Suppose there are $n$ bidders, each bidder $i$ with type space $\Theta_{i}$. Take a given direct-revelation mechanism. For each reported type profile $\theta=(\theta_{1},\dots,\theta_{n})$, suppose the allocation rule is that bidder $i$ wins with probability $p_{i}^{\theta}$. Define the interim allocation rule to be the mapping $\pi:\bigsqcup_{i}\Theta_{i}\to[0,1]$ where $\pi(\theta_{i})$ denotes the probability that bidder $i$ will win the item conditioned on reporting $\theta_{i}$ assuming others are bidding truthfully. Now suppose instead of taking a given a direct-revelation mechanism and defining the interim allocation rule, we start with an arbitrary function $\pi:\bigsqcup_{i}\Theta_{i}\to[0,1]$. Say that $\pi$ is a feasible interim allocation rule if $\pi$ can arise as an interim allocation rule from an actual direct-revelation mechanism’s allocation rule. Obviously, some functions $\pi$ are not feasible interim allocation rules. For example, $\pi\equiv 1$ is not feasible if $n\geq 2$ because no mechanism can guarantee everyone a probability $1$ of winning regardless of reported type profile.
Border’s Theorem gives a succinct characterization of which functions $\pi$ are feasible interim allocation rules. We state it below. It was proven by a sequence of previous work [border1991implementation, border2007reduced, mierendorff2011asymmetric].
Theorem C.1 (Border’s Theorem).
A function $\pi:\bigsqcup_{i}\Theta_{i}\to[0,1]$ is a feasible interim allocation rule if and only if for every measurable $\mathcal{T}=(\mathcal{T}_{1},\dots,\mathcal{T}_{n})\subseteq\Theta_{1}\times\dots\times\Theta_{n}$,
[…]
The problem of determining which probabilities $p_{i}\in[0,1]$ are interim allocation probabilities induced by allocation probabilities $p_{i}^{S}$ can be seen as a special case of Border’s Theorem. Specifically, assume on a given round in our mechanism that each agent $i$ bids with probability $\beta_{i}$, and we want to set allocation probabilities $p_{i}^{S}$ such that each agent’s probability of winning the item is $p_{i}$ conditioned on bidding. We can think of this problem as a special case of determining whether an interim allocation rule $\pi$ is feasible as follows. Think of each agent $i$ as a bidder with $2$ types: “bidding” and “not bidding” where the probability they have the “bidding” type is $\beta_{i}$. Set $\pi$ to be the function that is 00 on the “not bidding” type and $p_{i}$ on the “bidding” type for agent $i$. Determining whether it is feasible to guarantee each agent probability $p_{i}$ of winning conditioned on bidding is then equivalent to $\pi$ being a feasible interim allocation rule. Using Border’s Theorem to characterize whether $\pi$ is feasible, combined with the additional requirements that we never allocate the item to an agent who is not bidding and we always allocate the item to some agent if at least one agent bids, we obtain the following theorem (a generalization of Theorem 5.2 and a restatement of Theorem A.3).
Theorem C.2.
At a given round in our mechanism Generalized Budgeted Border, suppose each agent $i$ bids with probability $\beta_{i}$ independently across agents. Given arbitrary probabilities $p_{i}\in[0,1]$, there exist allocation probabilities $p_{i}^{S}$ such that the probability that $i$ wins the item conditioned on bidding is $p_{i}$ (i.e., the allocation probabilities $p_{i}^{S}$ induce the interim allocation probabilities $p_{i}$) if and only if
$\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$ (10)
and for every $I\subseteq[n]$,
$\sum_{i\in I}p_{i}\beta_{i}\leq 1-\prod_{i\in I}(1-\beta_{i}).$ (11)
We can see that the “only if” direction is obvious as follows. The left-hand side of (10) is the probability that some agent wins the item. The right-hand side of (10) is the probability that some agent bids. These must be equal since we allocate the item whenever at least one agent bids. Now consider (11). The left-hand side is the probability that an agent $i\in I$ wins the item. The right-hand side is the probability that $S$ contains an agent $i\in I$. Since an agent can only win the item if she bids, the left-hand side can be at most the right-hand side.
Now let us formally use Border’s Theorem to prove the “if” direction.
Proof of Theorem A.3.
Assume (10) and (11) hold. We do a reduction to the problem of determining feasibility of an interim allocation rule. Define a type space $\Theta_{i}=\{0_{i},1_{i}\}$ for each agent $i$ where $\Pr_{\theta_{i}\sim\Theta_{i}}(\theta_{i}=1_{i})=\beta_{i}$. Define $\pi$ by
$\pi(0_{i})$ $=0,\,\pi(1_{i})=p_{i}.$
Take any $(\mathcal{T}_{1},\dots,\mathcal{T}_{n})\subseteq\Theta_{1}\times\dots\times\Theta_{n}$. Without loss of generality, assume each $\mathcal{T}_{i}\subsetneq\Theta_{i}$. Let $I=\{i:\mathcal{T}_{i}=\{1_{i}\}\}$. Then,
[…]
by (11). By Theorem C.1, $\pi$ is a feasible interim allocation rule. Find a direct-revelation mechanism with interim allocation rule $\pi$ where $p_{i}^{\theta}$ is the probability that bidder $i$ wins if the reported type profile is $\theta$. Let $p_{i}^{S}=p_{i}^{\theta}$ where $\theta$ has $S=\{i\in[n]:\theta_{i}=1\}$. Notice that $0=\pi(0_{i})$ means that we never allocate the item to an bidder with type $0_{i}$, which implies that $p_{i}^{S}=0$ for $i\notin S$. For $i\in S$, $p_{i}=\pi(1_{i})$ is the probability that we allocate the item to a bidder with type $1_{i}$ conditioned on them having type $1_{i}$. By construction, this is the same conditional probability that we allocate the item to an agent $i$ conditioned on them bidding in Generalized Budgeted Border with allocation probabilities $p_{i}^{S}$. The sum $\sum_{i\in[n]}p_{i}\beta_{i}$ is the probability that bidder $i$ wins the item, and $1-\prod_{i\in[n]}(1-\beta_{i})$ is the probability that some agent $i$ has type $1_{i}$. By (10), the direct-revelation mechanism must allocate the item with probability $1$ to some agent if there is an agent with reported type $1_{i}$. Since $p_{i}^{S}=0$ unless $i\in S$, (10) implies that $\sum_{i\in S}p_{i}^{S}=1$, so $(p_{i}^{S})_{i\in S}$ is indeed a probability distribution over $i\in S$. ∎
C.2 A DMMF Proof of Theorem C.2
Although Theorem C.2 is a special case of Border’s Theorem, to aid intuition, we give novel proof of Theorem C.2 for our special case of allocation to bidding agents by simulating a different request-based allocation mechanism for the same setting.
We use the Dynamic Max-Min Fairness Mechanism (DMMF) introduced for our fair allocation setting by [fikioris2023online]. The DMMF mechanism goes as follows. In each round $t$, each agent $i$ can decide to request or not. Letting $S[t]$ be the set of agents that request, and $W_{i}[t]$ be the number of items that agent $i$ has won prior to and including round $t$, the principal allocates the item to the bidding agent $i\in S[t]$ that has the smallest value of $\frac{W_{i}[t-1]}{\alpha_{i}}$, the number of wins so far normalized by fair share.
The following was implicitly proven by [chido] via a Foster-Lyapunov argument.
Theorem C.3.
Define $Y_{i}[t]=\frac{W_{i}[t]}{\alpha_{i}}-\sum_{j=1}^{n}W_{j}[t]$. The vectors $Y[t]=(Y_{j}[t])_{j=1}^{n}$ form an irreducible Markov chain. Suppose each agent $i$ is bidding i.i.d. $\mathrm{Bernoulli}(\beta_{i})$ across rounds and independently across agents such that for any $\emptyset\subsetneq I\subsetneq[n]$,
$\frac{1-\prod_{i\in I}(1-\beta_{i})}{\sum_{i\in I}\alpha_{i}}>1-\prod_{j=1}^{n}(1-\beta_{j}).$ (12)
Then, the Markov chain $(Y[t])$ is positive recurrent, and
$\frac{W_{i}[T]}{T}\overset{\mathrm{a.s.}}{\to}\alpha_{i}\left(1-\prod_{j=1}^{n}(1-\beta_{j})\right).$ (13)
Now we give our DMMF proof of Border’s Theorem, which will roughly go as follows. Given interim allocation probabilities $p_{i}$ satisfying the conditions of Theorem A.3, we shall define fair shares $\alpha_{i}$ such that when we run DMMF, the long-run fraction of items that agent $i$ wins is $\beta_{i}p_{i}$. Let $\mu$ be the distribution over $S\subseteq[n]$ where each $i\in S$ independently with probability $\beta_{i}$. We shall show that these fraction items come from the expectation of $\beta_{i}p_{i}^{S}$ over sets $S\sim\mu$ of bidding agents, where $p_{i}^{S}$ is the long-run fraction of times $t$ that agent $i$ would have won had they requested and $S$ were the set of agents who requested at time $t$. These allocation probabilities $p_{i}^{S}$’s will therefore induce the interim allocation probabilities $p_{i}$.
DMMF proof of Theorem C.2.
Necessity of Border’s Criterion for the $p_{i}^{S}$’s to exist was previously argued to be obvious in the remarks after Theorem C.2. For sufficiency, let $(\beta_{i})$ be bid rates and $(p_{i})$ be interim allocation probabilities satisfying Border’s Criterion. We shall assume that (11) holds with strict inequality for $I\subsetneq[n]$; a routine topological argument given in Appendix D extends our proof to the general case.
Define fair shares
$\alpha_{i}=\frac{\beta_{i}p_{i}}{1-\prod_{j=1}^{n}(1-\beta_{j})}.$
to be used in DMMF. By (10), the fair shares indeed satisfy $\sum_{i=1}^{n}\alpha_{i}=1$. With these fair shares, by rearrangement, (11) holding with strict inequality and (12) are equivalent. Therefore, the irreducible Markov chain $(Y[t])$ in Theorem C.3 is positive recurrent.
In the DMMF process, let $\mathcal{S}_{i}[t]$ be the set of subsets $S\subseteq[n]$ of agents such that if $S$ were to be the set of agents who requested at time $t$, agent $i$ would have won the item,
$\mathcal{S}_{i}[t]=\left\{S\subseteq[n]:i=\operatorname*{arg\,min}_{j\in S}\frac{W_{j}[t-1]}{\alpha_{j}}\right\}.$
Observe that $\mathcal{S}_{i}[t]$ can be written as a function of $Y[t-1]$. By the pointwise ergodic theorem, for any $S\subseteq[n]$,
$\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{1}\{S\in\mathcal{S}_{i}[t]\}\overset{\mathrm{a.s.}}{\to}p_{i}^{S}$
for some constant $p_{i}^{S}\in[0,1]$.
Let $S[t]$ be the actual set of agents that request at time $t$. By (13) and the fact that agent $i$ actually wins the item at time $t$ if and only if $S[t]\in\mathcal{S}_{i}[t]$,
$\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{1}\{S[t]\in\mathcal{S}_{i}[t]\}\overset{\mathrm{a.s.}}{\to}\alpha_{i}\left(1-\prod_{j=1}^{n}(1-\beta_{j})\right).$
Recall that $\mu$ is the distribution over $S\subseteq[n]$ where each $i\in S$ independently with probability $\beta_{i}$. Suppose we sample the set $S\sim\mu$. For each $t$, observe that both $S$ and $S[t]$ have distribution $\mu$ and both are independent of $\mathcal{S}_{i}[t]$. Therefore,
$\frac{1}{T}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}_{S\sim\mu}[\boldsymbol{1}\{S\in\mathcal{S}_{i}[t]\}]=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\boldsymbol{1}\{S[t]\in\mathcal{S}_{i}[t]\}].$ (14)
Taking $T\to\infty$ and applying bounded convergence, we obtain
$\operatorname*{\mathbb{E}}_{S\sim\mu}[p_{i}^{S}]=\alpha_{i}\left(1-\prod_{j=1}^{n}(1-\beta_{j})\right)=\beta_{i}p_{i}.$ (15)
We observe that $p_{i}^{S}=0$ if $i\notin S$, and so by definition, the $p_{i}^{S}$’s induce the interim allocation probabilities $p_{i}$. ∎
C.3 A Flow Network Proof of Theorem C.2
A proof of Border’s Theorem based on flow networks and the max-flow min-cut theorem was discovered by [che2013generalized]. We will give flow network proof in the special case of Border’s Theorem that we use (Theorem C.2) here for convenience.
[…]
Figure 4: Flow network that can be used to prove Theorem A.3. We let $S^{\prime}$ be the random set of bidding agents where each agent $i$ lies in $S^{\prime}$ independently with probability $\beta_{i}$. Then, there are three kind of edges: edges whose flow corresponds to the probability of observing a specific $S^{\prime}$ (left), edges whose flow corresponds to how we randomly allocate the item condition on observing a specific $S^{\prime}$ (middle), and edges whose flow represent the probability that a specific agent gets the item (right). There is a flow of value $\Pr(S^{\prime}\neq\emptyset)$ if and only if there exist allocation probabilities $p_{i}^{S}$ inducing the interim allocation probabilities $p_{i}$. In other words, the flows $p_{i}^{S}\Pr(S^{\prime}=S)$ in the middle transform the probabilities that agents in a certain set $S$ bid to an agent $i\in S$ being allocated. We obtain the conditions in Theorem A.3 by analyzing every minimum-cut of this network.
Flow network proof of Theorem C.2.
As argued in the remarks after Theorem C.2, (10) is necessary for the $p_{i}^{S}$’s to exist, so we assume it.
Let $\mu$ be the distribution over $S^{\prime}\subseteq[n]$ where each $i\in S^{\prime}$ independently with probability $\beta_{i}$. Create an $s$-$t$ flow network as follows. For each nonempty $S\subseteq[n]$ create a node $u_{S}$ and an edge $(s,u_{S})$ with capacity
$c(s,u_{S})=\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S).$ (16)
For each $i\in[n]$, create a node $v_{i}$. For each $S$ such that $i\in S$, add an edge $(u_{S},v_{i})$ with infinite capacity. Also, add an edge $(v_{i},t)$ with capacity
$c(v_{i},t)=p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime}).$ (17)
The flow network is depicted in Fig. 4.
The cut with $s$ on one side and everything else on the other has capacity
$\sum_{S\subseteq[n]:S\neq\emptyset}c(s,u_{S})=\sum_{S\subseteq[n]:S\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\neq\emptyset)=1-\prod_{i\in S}(1-\beta_{i}).$ (18)
The cut with $t$ on one side and everything else on the other has capacity
$\sum_{i\in[n]}c(v_{i},t)=\sum_{i\in[n]}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})=\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i}),$ (19)
using (10) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\beta_{i})$, it suffices to show that (11) holds only if there is a feasible flow of flow value equal to this cut capacity.
Take any minimum-capacity $s$-$t$ cut $(A,B)$. Since the edges $(u_{S},v_{i})$ have infinite capacity, if $v_{i}\in B$ then $u_{S}\in B$ for any $S$ such that $i\in S$. Conversely, for any $S$, if $v_{i}\in A$ for every $i\in S$, then $u_{S}\in A$ since there are no edges coming out of $u_{S}$ except the $(u_{S},v_{i})$. Thus, the cut $(A,B)$ is completely characterized by which nodes $v_{i}\in B$. Let $I\subseteq[n]$ be the subset such that $v_{i}\in B$ if and only if $i\in I$. The total capacity of this cut is
$\begin{split}c(A,B)&=\sum_{S\subseteq[n]:S\cap I\neq\emptyset}c(s,u_{S})+\sum_{i\notin I}c(v_{i},t)\\ &=\sum_{S\subseteq[n]:S\cap I\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)+\sum_{i\notin I}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})\\ &=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\cap I\neq\emptyset)+\sum_{i\notin I}p_{i}\beta_{i}\\ &=1-\prod_{i\in I}(1-\beta_{i})+1-\prod_{i\in[n]}(1-\beta_{i})-\sum_{i\in I}p_{i}\beta_{i},\end{split}$ (20)
using (10) for the last line. Rearranging, one can see that (11) is equivalent to the above being least $1-\prod_{i\in[n]}(1-\beta_{i})$. ∎
C.4 Theorem C.2 with Arbitrary Upper Bounds
The proof similar to the proof of Border’s Theorem given in [che2013generalized]. We construct a flow network that is feasible if and only if there are allocation probabilities inducing given interim allocation probabilities that satisfy some upper bounds. We obtain inequalities that determine whether the flow network is feasible by analyzing every possible minimum cut.
Theorem C.4.
Given upper bounds $\bar{p}_{i}^{S}$, there exists $p_{i}^{S}$’s that induce interim allocation probabilities $p_{i}$ such that $p_{i}^{S}\leq\bar{p}_{i}^{S}$ for every $S\neq\emptyset$ if and only if
$\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$ (21)
and for any $I\subseteq[n]$ and $\mathcal{S}\subseteq\{S\subseteq[n]:S\cap I\neq\emptyset\}$,
$\sum_{i\in I}p_{i}\beta_{i}+\prod_{i\in I}(1-\beta_{i})+\sum_{S\in\mathcal{S}}\left(\prod_{i\in S}\beta_{i}\right)\left(\prod_{i\notin S}(1-\beta_{i})\right)\left(1-\sum_{i\in S\cap I}\bar{p}_{i}^{S}\right)\leq 1.$ (22)
Proof.
Since (21) is necessary for the $p_{i}^{S}$’s to exist in Theorem 5.2, it is also necessary here, so we assume it.
Create an $s$-$t$ flow network as follows. Let $\mu$ be the probability distribution over subsets $S^{\prime}\subseteq[n]$ where each $i\in S^{\prime}$ independently with probability $\beta_{i}$. For each nonempty $S\subseteq[n]$ create a node $u_{S}$ and connect it with an edge $(s,u_{S})$ to the source node with capacity
$c(s,u_{S})=\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S).$
For each $i$, create a node $v_{i}$. For each $S\neq\emptyset$ such that $i\in S$, add an edge $(u_{S},v_{i})$ with capacity
$c(u_{S},v_{i})=\bar{p}_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$
Also, add an edge $(v_{i},t)$ with capacity
$c(v_{i},t)=p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime}).$
The flow network is depicted in Fig. 5.
[…]
Figure 5: Flow network that is used in the proof of Theorem C.4. The flow network is the same as Fig. 4 except the middle edges $(u_{S},v_{i})$ have capacity $\bar{p}_{i}^{S}\Pr(S^{\prime}=S)$ to enforce the upper bounds on the allocation probabilities.
The cut with $s$ on one side and everything else on the other has capacity
$\sum_{S\subseteq[n]:S\neq\emptyset}c(s,u_{S})=\sum_{S\subseteq[n]:S\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\neq\emptyset)=1-\prod_{i\in S}(1-\beta_{i}).$
The cut with $t$ on one side and everything else on the other has capacity
$\sum_{i\in[n]}c(v_{i},t)=\sum_{i\in[n]}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})=\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i}),$
using (21) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ satisfying upper bounds $p_{i}^{S}\leq\bar{p}_{i}^{S}$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\beta_{i})$, it suffices to show that (22) holds only if there is a feasible flow of flow value equal to this cut capacity.
Take any minimum-capacity $s$-$t$ cut $(A,B)$. For any $S$, if $v_{i}\in A$ for every $i\in S$, we can assume $u_{S}\in A$ since there are no edges coming out of $u_{S}$ except the $(u_{S},v_{i})$. Thus, the cut $(A,B)$ is completely characterized by which nodes $v_{i}\in B$ and which nodes $u_{S}\in A$ for $S$ containing some $i$ such that $v_{i}\in B$. Let $I=\{i\in[n]:v_{i}\in B\}$ and $\mathcal{S}=\{S\subseteq 2^{[n]}:S\cap I\neq\emptyset,u_{S}\in A\}$. The total capacity of this cut is
$\begin{split}c(A,B)&=\sum_{S\notin\mathcal{S}:S\cap I\neq\emptyset}c(s,u_{S})+\sum_{S\in\mathcal{S}}\sum_{i\in S\cap I}c(u_{S},v_{i})+\sum_{i\notin I}c(v_{i},t)\\ &=\sum_{S:S\cap I\neq\emptyset}c(s,u_{S})-\sum_{S\in\mathcal{S}}c(s,u_{S})+\sum_{S\in\mathcal{S}}\sum_{i\in S\cap I}c(u_{S},v_{i})+\sum_{i\notin I}c(v_{i},t)\\ &=\sum_{S:S\cap I\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)-\sum_{S\in\mathcal{S}}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)+\sum_{S\in\mathcal{S}}\sum_{i\in S\cap I}\bar{p}_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)+\sum_{i\notin I}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})\\ &=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\cap I\neq\emptyset)-\sum_{S\in\mathcal{S}}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)\left(1-\sum_{i\in S\cap I}\bar{p}_{i}^{S}\right)+\sum_{i\notin I}p_{i}\beta_{i}\\ &=1-\prod_{i\in I}(1-\beta_{i})-\sum_{S\in\mathcal{S}}\left(\prod_{i\in S}\beta_{i}\right)\left(\prod_{i\notin\mathcal{S}}(1-\beta_{i})\right)\left(1-\sum_{i\in S\cap I}\bar{p}_{i}^{S}\right)\\ &\quad+1-\prod_{i\in[n]}(1-\beta_{i})-\sum_{i\in I}p_{i}\beta_{i}\end{split}$
using (21) for the last line. Rearranging, (22) is equivalent to the above being least $1-\prod_{i\in[n]}(1-\beta_{i})$. ∎
Appendix D Deferred Proofs
D.1 Deferred Proofs of Corollaries A.4 and A.5
The proof of Corollary A.4 is almost identical to proof of Theorem 4.2 given in Section 5.2.
Proof of Corollary A.4.
We prove that the choice of $p_{i}=\frac{1-\prod_{j=1}^{n}(1-\gamma\alpha_{j})}{\gamma}$ satisfies the conditions of Theorem A.3. Observe that (8) holds with this choice of $p_{i}$.
Also, observe that (3) holds for $I=\emptyset$. To show that (3) holds for $I\neq\emptyset$, define $a_{I}$ as
$a_{I}=\frac{1-\prod_{i\in I}(1-\beta_{i})}{\sum_{i\in I}\beta_{i}}.$
With this choice of $p_{i}$, (3) says
$\sum_{i\in I}p_{i}\beta_{i}=\frac{1-\prod_{j=1}^{n}(1-\beta_{j})}{\gamma}\sum_{i\in I}\beta_{i}\leq 1-\prod_{i\in I}(1-\beta_{i}).$
This holds if and only if $a_{I}\geq\frac{1-\prod_{j=1}^{n}(1-\beta_{j})}{\gamma}$. It suffices to show $a_{I}$ is nonincreasing in $I$ since $a_{[n]}=\frac{1-\prod_{j=1}^{n}(1-\beta_{j})}{\gamma}$. To do this, we can simply follow the same proof that $a_{I}$ as defined in the proof of Theorem 4.2 is nonincreasing, but replace each $\alpha_{i}$ with $\beta_{i}$. We did not use the fact that $\sum_{i=1}^{n}\alpha_{i}=1$ for that part of the proof of Theorem 4.2, so everything still works. ∎
The proof of Corollary A.5 is an application of Corollary A.4.
Proof of Corollary A.5.
With a $\mathrm{Uniform}([0,1])$ value distribution, the $\beta$-ideal utility is
$v^{\star}(\beta)=\operatorname*{\mathbb{E}}_{V\sim\mathrm{Uniform}([0,1])}[V\boldsymbol{1}\{V>1-\beta\}]=\frac{1}{2}\beta(2-\beta).$
If $\beta=\gamma\alpha$, then
$v^{\star}(\beta)=\frac{\gamma(1-\gamma\alpha)}{2-\alpha}v^{\star}(\alpha).$
Using Corollary A.4, we can find interim allocation probabilities $p_{i}$ such that,
$p_{i}v^{\star}(\beta)\geq\frac{1-e^{-\gamma}}{\gamma}v^{\star}(\beta)\geq\frac{(1-e^{-\gamma})(2-\gamma\alpha)}{2-\alpha}.$
The result follows from substituting $\alpha=\frac{1}{n}$ and $\gamma=\Theta(\log n)$. ∎
D.2 Deferred Proofs of Lemma 5.1, Propositions 4.1 and A.1, Theorems 4.3 and A.2
We shall prove these in the context of our more general mechanism Generalized Budgeted Border, which is the same as Budgeted Robust Border except the each agent gets $\beta_{i}(1+\delta_{i}^{T})$ budget of bid tokens where $\beta_{i}\in[0,1]$ is a parameter. Budgeted Robust Border is Generalized Budgeted Border when each $\beta_{i}=\alpha_{i}$. See Appendix A for details.
To prove results about agents’ behavior in the mechanism, it will be useful to use the following imaginary game. This imaginary game will be the same game, but we do not enforce budgets and allow agents to bid regardless of whether they have budget remaining. Let $\tilde{b}_{i}^{t}$, $\tilde{U}_{i}[t]$, and $\tilde{i}^{t}$ be the bids, utilities, and winners in this imaginary game, respectively. We couple the imaginary game and the actual game such that the agents have the same values $V_{i}[t]$, and $\tilde{b}_{i}^{t}=b_{i}^{t}$, $\tilde{U}_{i}[t]=U_{i}[t]$, and $\tilde{i}^{t}=i^{t}$ at all times $t$ in which all agents have budget remaining.
We first prove a lemma about the best strategy for an agent $i$ assuming they 1) do not want to exceed their budget in expectation and 2) the other agents $j\neq i$ are acting in a specific way.
Lemma D.1.
Fix an agent $i$ and assume all other agents $j\neq i$ are playing in a way such that the sets $S_{\neq i}[t]=\{j\neq i:b_{j}^{t}=1\}$ of bidding agents $j\neq i$ are i.i.d. across rounds drawn from some distribution $\nu$ over subsets of $[n]\setminus\{i\}$. Suppose agent $i$ is trying to maximize her imaginary expected utility subject to the constraint that she does not exceed her budget in expectation; that is, she is choosing $(\tilde{b}_{i}^{t})$ to solve
$\max\,\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]\quad\text{subject to}\quad\sum_{t=1}^{T}\mathbb{E}[\tilde{b}_{i}^{t}]\leq\beta_{i}^{\prime}T$ (23)
where $\beta_{i}^{\prime}=\beta_{i}(1+\delta_{i}^{T})$. Her optimal strategy is to choose $(\tilde{b}_{i}^{t})$ to be a $\beta_{i}^{\prime}$-aggressive strategy. Letting
$p_{i}(\nu)=\sum_{S\subseteq[n]}p_{i}^{S}\Pr_{S_{\neq i}\sim\nu}(S_{\neq i}\cup\{i\}=S)$ (24)
be the probability that agent $i$ wins a round conditioned on bidding, the $\beta_{i}^{\prime}$-aggressive strategy yields agent $i$ utility
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]=p_{i}(\nu)v^{\star}(\beta_{i}^{\prime}).$
Notice that Lemma 5.1 is a special case of the above lemma with parameters $\beta_{i}=\alpha_{i}$, $\delta_{i}^{T}=0$, and $\nu$ being the distribution over subsets $S_{\neq i}$ where each $j\in S_{\neq i}$ independently with probability $\alpha_{j}$. In the proof, we use the notion of $\beta$-ideal utility and $\beta$-ideal utility probability function from Definition A.1 defined in Appendix A.
Proof of Lemma D.1.
For any strategy $(\tilde{b}_{i}^{t})$, at any time $t$,
$\begin{split}\mathbb{E}[\tilde{U}_{i}[t]]=\mathbb{E}[V_{i}[t]\boldsymbol{1}\{i=\tilde{i}^{t}\}]=\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}\boldsymbol{1}\{i=\tilde{i}^{t}\}]=\mathbb{E}[\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}\boldsymbol{1}\{i=\tilde{i}^{t}\}\mid\tilde{b}_{i}^{t}]]\\ =\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]\mathbb{E}[\boldsymbol{1}\{i=\tilde{i}^{t}\}\mid\tilde{b}_{i}^{t}=1]=p_{i}(\nu)\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}].\end{split}$ (25)
Thus, agent $i$’s maximization problem is equivalent to maximizing $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]$ subject to the same constraint in (23). It is clear for any feasible solution $(\tilde{b}_{i}^{t})$ that the solution that bids at time $t$ with probability $\rho[t](V_{i}[t])=\Pr(\tilde{b}_{i}^{t}=1\mid V_{i}[t])$ is also a feasible solution with the same objective value. Thus, we can rewrite the agent’s maximization problem in terms of maximizing over $\rho[t](V_{i}[t])$, i.e., the following maximization problem over measurable functions $\rho[t]:[0,\infty)\to[0,1]$:
$\frac{1}{T}\max\,\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]\rho[t](V_{i}[t])]\quad\text{subject to}\quad\sum_{t=1}^{T}\mathbb{E}[\rho[t](V_{i}[t])]\leq\beta_{i}^{\prime}T.$
Given any optimal solution $(\rho[t])$ to the above, observe that setting $\rho^{\star}[t]=\frac{1}{T}\sum_{s=1}^{T}\rho[s]$ is also a feasible solution with the same objective value. Observe then that $\mathbb{E}[\rho^{\star}[t](V_{i}[t])]=\beta_{i}^{\prime}$, and so $\rho^{\star}[t]$ is a feasible solution in (A.1), the definition of $\beta_{i}^{\prime}$-ideal utility, and it must maximize the same objective $\mathbb{E}[V_{i}[t]\rho[t](V_{i}[t])]]$. Therefore, $\rho^{\star}[t]$ is exactly the $\beta_{i}^{\prime}$-ideal utility probability function $(\rho_{i}^{\beta_{i}^{\prime}})^{\star}$, so the optimal bidding strategy $(\tilde{b}_{i}^{t})^{\star}$ to solve (23) is precisely a $\beta_{i}^{\prime}$-aggressive strategy.
By (25), under such a $\beta_{i}^{\prime}$-aggressive strategy, agent $i$ obtains utility $p_{i}(\nu)v^{\star}(\beta_{i}^{\prime})$. ∎
To start relating the imaginary game to the actual game, we use Chernoff bounds to show that agents $i$ who use $\beta_{i}$-aggressive strategies will not run out of budget with high probability.
Lemma D.2.
If agent $i$ uses a $\beta_{i}$-aggressive strategy, the probability that they run out of budget is at most $O\left(\frac{1}{T^{2}}\right)$.
Proof.
By the Chernoff bound,
$\Pr\left(\sum_{t=1}^{T}b_{i}^{t}\geq\beta_{i}(1+\delta_{i}^{T})T\right)\leq\exp\left(-\frac{(\delta_{i}^{T})^{2}\beta_{i}T}{2+\delta_{i}^{T}}\right).$
The result follows from substituting $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$. ∎
We shall use the following lemma to obtain high probability bounds on the agents’ utilities in the actual game. Remember that we coupled in the imaginary game and the actual game such that the agents have the same values $V_{i}[t]$, and $\tilde{b}_{i}^{t}=b_{i}^{t}$, $\tilde{U}_{i}[t]=U_{i}[t]$, and $\tilde{i}^{t}=i^{t}$ at all times $t$ in which all agents have budget remaining. This implies that the strategy used by a player $i$ in the imaginary game directly translates to a strategy used by $i$ in the actual game. We define a $\beta$-aggressive strategy in the imaginary game to be one in which agent $i$ bids when her value is in the top $\beta$-quantile of her value distribution. Notice that if player $i$ is playing a $\beta$-aggressive strategy in the imaginary game, then she is playing a $\beta$-aggressive strategy in the actual game (since the only difference is the budget constraint).
Lemma D.3.
Fix an agent $i$ playing a $\beta_{i}$-aggressive strategy. Suppose the other agents $j\neq i$ are playing in the imaginary game with strategies as in Lemma D.1. Let $E_{1}$ be an event on which agents $j\neq i$ do not run out of budget in the actual game. Then, on an subevent of $E_{1}$ with probability at least $\Pr(E_{1})-O(1/T^{2})$,
$\left|\frac{1}{T}\sum_{t=1}^{T}\tilde{U}_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})\right|\leq O\left(\sqrt{\frac{\log T}{T}}\right).$
Proof.
Observe that the random variables $\tilde{U}_{i}[t]=V_{i}[t]\boldsymbol{1}\{\tilde{i}^{t}=i\}$ are i.i.d. By (25), each has mean $p_{i}(\nu)\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]$. Since player $i$ is playing a $\beta_{i}$-aggressive strategy, this mean is $\mathbb{E}[\tilde{U}_{i}[t]]=p_{i}(\nu)v_{i}^{\star}(\beta_{i})$. Recall that we assume the value distribution $\mathcal{F}_{i}$ is bounded so the $U_{i}$ are bounded by some $\bar{v}$. Let $\epsilon>0$. By Hoeffding’s inequality,
$\Pr\left(\left|\sum_{t=1}^{T}\tilde{U}_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})T\right|\geq\epsilon\right)\leq 2\exp\left(-\frac{2\epsilon^{2}}{\bar{v}^{2}T}\right)$ (26)
Let $E_{2}$ be the event that the above event does not occur. Let $E_{3}$ be the event that agent $i$ does not run out of budget. Let $E=E_{1}\cap E_{2}\cap E_{3}$. On $E$,
$\left|\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})\right|=\left|\frac{1}{T}\sum_{t=1}^{T}\tilde{U}_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})\right|\leq\frac{\epsilon}{T}.$
Substituting $\epsilon=\bar{v}\sqrt{T\ln T}$, the above is at most $O\left(\sqrt{\frac{\log T}{T}}\right)$, and by also substituting this $\epsilon$ into (26) and by Lemma D.2, we have $\Pr(E)\geq\Pr(E_{1})-O(1/T^{2})$, giving the result. ∎
The following lemma gives some form of continuity in the ideal utility that we need to bound the utility an agent can obtain from deviating from the proposed equilibrium.
Lemma D.4.
Let $\beta_{i}^{\prime}=(1+\delta_{i}^{T})\beta_{i}$. Then,
$v_{i}^{\star}(\beta^{\prime})-v_{i}^{\star}(\beta)\leq\delta_{i}^{T}v_{i}^{\star}(\beta_{i}).$
Proof.
It was proven in [fikioris2023online] that $\beta\mapsto v_{i}^{\star}(\beta)$ is concave. The lemma statement follows from concavity and the fact that $v_{i}^{\star}(0)=0$. ∎
With those lemmas, we can now prove that each player following a $\beta_{i}$-aggressive strategy and give their utility guarantee. We use $p_{i}$ to denote player $i$’s interim allocation probability, which we define in Definition A.2. Equivalently, $p_{i}$ denotes $p_{i}(\nu)$ as defined in (24) when $\nu$ is the distribution over $S_{\neq i}$ where $j\in S_{\neq i}$ independently with probability $\beta_{j}$. The below theorem is Propositions A.1 and A.2 combined. It is a direct generalization of Proposition 4.1, which can be obtained by setting $\beta_{i}=\alpha_{i}$ for each $i$. When using $\beta_{i}=\alpha_{i}$ and Lemmas 6.2 and 6.3 to set the allocation probabilities, we also obtain Theorem 4.3.
Theorem D.5.
Suppose we run Generalized Budgeted Border with slack parameters $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$. Each player $i$ playing a $\beta_{i}$-aggressive is an $O\left(\sqrt{\frac{\log T}{T}}\right)$-approximate Nash equilibrium. At this approximate equilibrium, with probability at least $1-O(1/T^{2})$, player $i$ gets utility
$\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq p_{i}v^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log T}{T}}\right).$
Proof.
Suppose every agent $i$ is using a $\beta_{i}$-aggressive strategy. Let $E_{1}$ be the event that no agent runs out of budget. By Lemma D.2, $\Pr(E_{1})\geq 1-O(1/T^{2})$. Using Lemma D.3, there is an event $E$ of probability at least $1-O(1/T^{2})$ on which
$\frac{1}{T}\sum_{t=1}^{T}\tilde{U}_{i}[t]\geq p_{i}v^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log T}{T}}\right).$
This establishes the high probability utility guarantee if every agent $i$ is playing a $\beta_{i}$-aggressive strategy. Now let us show this strategy profile is indeed a Nash equilibrium. The high probability utility guarantee translates to a utility guarantee in expectation in that
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]\geq p_{i}v^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log T}{T}}\right).$ (27)
Now we upper bound the utility of any deviating strategy by player $i$, still assuming players $j\neq i$ are following a $\beta_{j}$-aggressive strategy. Any strategy $(b_{i}^{t})$ used by player $i$ in the actual game satisfies the budget constraint $\sum_{t=1}^{T}b_{i}^{T}\leq\beta_{i}(1+\delta_{i}^{T})T$ almost surely. In particular, it satisfies the budget constraint in expectation, so we can use Lemma 5.1 to conclude that under the strategy $(b_{i}^{t})$ in the imaginary game,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]\leq p_{i}v^{\star}(\beta_{i}(1+\delta_{i}^{T})).$
Therefore,
$\begin{split}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]&=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]\boldsymbol{1}_{E_{1}}]+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]\boldsymbol{1}_{E_{1}^{c}}]\\ &\leq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]\boldsymbol{1}_{E_{1}^{c}}]\\ &\leq p_{i}v_{i}^{\star}(\beta_{i}(1+\delta_{i}^{T}))+O\left(\frac{1}{T^{2}}\right).\end{split}$ (28)
By (27) and (28), by deviating from a $\beta_{i}$-aggressive strategy, player $i$ can only gain an additive utility difference of
$p_{i}v_{i}^{\star}(\beta_{i}(1+\delta_{i}^{T}))T-p_{i}v^{\star}(\beta_{i})T+O\left(\sqrt{\frac{\log T}{T}}\right).$
By substituting $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$ and using Lemma D.4, this implies that this additive difference is at most $O\left(\sqrt{\frac{\log T}{T}}\right)$, thus proving the theorem.
D.3 Deferred Proofs of Theorems B.1, B.2, and B.3
In all the proofs in this subsection, we use the following notation. Let $\tilde{b}_{i}^{t}$ be the bids of agent $i$ corresponding to a $\beta_{i}$-aggressive strategy in the imaginary game introduced in Section D.2 where there is no budget constraint that agree with the actual bids $b_{i}^{t}$ at times $t$ where agent $i$ has budget. Let $U_{i}[t,T]$ denote the utility of player $i$ gained at time $t$ under the policy $\tilde{b}_{i}^{t}$ if only the $n$ constraints $\sum_{t=1}^{T}b_{k}^{t}\leq\beta_{k}(1+\delta_{k}^{T})T$ were enforced. If agent $i$ deviates to a policy $(b_{i}^{t})$, we let $U_{i}^{\prime}[t]$ and $U_{i}^{\prime}[t,T]$ be analogous to $U_{i}[t]$ and $U_{i}[t,T]$ for the deviating policy.
Proof of Theorem B.1.
By Lemma D.2, the event $E_{1}^{s}$ that no agent runs out of budget at time $s$ has probability at least $1-O(1/s^{2})$. By the union bound, the event $E_{1}=\bigcap_{s=\lceil\sqrt{t}\rceil}^{t}E_{1}^{s}$ that no agent runs out of budget at any time $\lceil\sqrt{t}\rceil$ and $t$ has probability at least $1-\sum_{s=\lceil\sqrt{t}\rceil}^{t}(1-\Pr(E_{1}^{s}))=1-O(1/\sqrt{t})$. By Lemma D.3, on an subevent $E$ of $E_{1}$ with probability at least $1-O(1/\sqrt{t})$,
$\frac{1}{t}\sum_{s=1}^{t}U_{i}[s,t]=\frac{1}{t}\sum_{s=1}^{t}U_{i}[s,t]\geq p_{i}v_{i}^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log t}{t}}\right).$
Since on $E_{1}$, we have $U_{i}[s,t]=U_{i}[s]$ for each $s$ between $\lceil\sqrt{t}\rceil$ and $t$, we obtain on $E$,
$\frac{1}{t}\sum_{s=1}^{t}U_{i}[s]\geq\frac{1}{t}\sum_{s=\lceil\sqrt{t}\rceil}^{t}U_{i}[s,t]\geq\frac{1}{t}\sum_{s=1}^{t}U_{i}[s,t]-\frac{1}{t}\sum_{s=1}^{\lceil\sqrt{t}\rceil-1}V_{i}[s]\geq p_{i}v_{i}^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log t}{t}}\right),$
establishing the utility guarantee.
Suppose agent $i$ deviates to a policy $(b_{i}^{t})^{\prime}$. By Theorem D.5,
$\begin{split}\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}^{\prime}[s]]&=\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}^{\prime}[s]\boldsymbol{1}_{E_{1}}]+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}^{\prime}[s]\boldsymbol{1}_{E_{1}^{c}}]\\ &=\frac{1}{t}\sum_{s=1}^{\lceil\sqrt{t}\rceil}\mathbb{E}[U_{i}^{\prime}[s]]+\frac{1}{t}\sum_{s=\lceil\sqrt{t}\rceil}^{t}\mathbb{E}[U_{i}^{\prime}[s,t]]+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}^{\prime}[s]\boldsymbol{1}_{E_{1}^{c}}]\\ &\leq\frac{1}{t}\sum_{s=1}^{\lceil\sqrt{t}\rceil-1}\mathbb{E}[U_{i}^{\prime}[s]]+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s,t]]+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}^{\prime}[s]\boldsymbol{1}_{E_{1}^{c}}]\\ &=\frac{1}{t}\sum_{s=1}^{\lceil\sqrt{t}\rceil-1}\mathbb{E}[U_{i}^{\prime}[s]]+\frac{1}{t}\sum_{s=1}^{\lceil\sqrt{t}\rceil-1}\mathbb{E}[U_{i}[s,t]\boldsymbol{1}_{E_{1}}]+\frac{1}{t}\sum_{s=\lceil\sqrt{t}\rceil}^{t}\mathbb{E}[U_{i}[s,t]\boldsymbol{1}_{E_{1}}]\\ &\quad+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s,t]\boldsymbol{1}_{E_{1}^{c}}]+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}^{\prime}[s]\boldsymbol{1}_{E_{1}^{c}}]\\ &=\frac{1}{t}\sum_{s=1}^{\lceil\sqrt{t}\rceil-1}\mathbb{E}[U_{i}^{\prime}[s]]+\frac{1}{t}\sum_{s=1}^{\lceil\sqrt{t}\rceil-1}\mathbb{E}[U_{i}[s,t]\boldsymbol{1}_{E_{1}}]+\frac{1}{t}\sum_{s=\lceil\sqrt{t}\rceil}^{t}\mathbb{E}[U_{i}[s]\boldsymbol{1}_{E_{1}}]\\ &\quad+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s,t]\boldsymbol{1}_{E_{1}^{c}}]+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}^{\prime}[s]\boldsymbol{1}_{E_{1}^{c}}]\\ &\leq\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s]]+O\left(\sqrt{\frac{\log t}{t}}\right).\end{split}$
By substituting $t=T$, we see that everyone playing a $\beta_{i}$-aggressive strategy is indeed an $O\left(\sqrt{\frac{\log T}{T}}\right)$-equilibrium. ∎
Proof of Theorem B.2.
Assume without loss of generality that the agents $j\neq i$ never bid when they’re out of budget. As in proof of Theorem B.1, the event $E_{1}$ that agent $i$ does not run out of budget at any time between $\lceil\sqrt{t}\rceil$ and $t$ has probability at least $1-O(1/\sqrt{t})$. Then, on $E_{1}$, no one runs out of budget between time $\lceil\sqrt{t}\rceil$ and $t$. Using Theorem 4.3, there is an subevent $E$ of $E_{1}$ of probability at least $1-O(1/\sqrt{t})$ on which
$\begin{split}\frac{1}{t}\sum_{s=1}^{t}U_{i}[s]&\geq\frac{1}{t}\sum_{s=\lceil\sqrt{t}\rceil}^{t}U_{i}[s,t]\\ &\geq\left(\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2}\right)v_{i}^{\star}-O\left(\sqrt{\frac{\log t}{t}}\right)-\frac{1}{t}\sum_{s=1}^{\lceil\sqrt{t}\rceil-1}V_{i}[s]\\ &\geq\left(\frac{1}{2}+\frac{1}{2}\alpha_{i}^{2}\right)v_{i}^{\star}-O\left(\sqrt{\frac{\log t}{t}}\right).\end{split}$
Proof of Theorem B.3.
By Lemma D.2 and the Borel-Cantelli Lemma, there exists a random time $t_{0}$ such that $\sum_{s=1}^{t}\tilde{b}_{i}^{s}\leq\beta_{i}(1+\delta_{i}^{t})t$ for all $t>t_{0}$ and all agents $i$ where $t_{0}<\infty$ almost surely. Suppose agent $i$ deviates to a policy $(b_{i}^{s})^{\prime}$. Using Theorem D.5,
$\begin{split}\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s]]&=\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t_{0}}U_{i}[s]\right]+\mathbb{E}\left[\frac{1}{t}\sum_{s=t_{0}+1}^{t}U_{i}^{\prime}[t,T]\right]\\ &\leq\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t_{0}}U_{i}[s]\right]+\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t}U_{i}^{\prime}[t,T]\right]\\ &\leq\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t_{0}}U_{i}[s]\right]+\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t}U_{i}[s,t]\right]+O\left(\sqrt{\frac{\log t}{t}}\right)\\ &=\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t_{0}}U_{i}[s]\right]+\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t_{0}}U_{i}[s,t]\right]+\mathbb{E}\left[\frac{1}{t}\sum_{s=t_{0}+1}^{t}U_{i}[s]\right]+O\left(\sqrt{\frac{\log t}{t}}\right)\\ &\leq 2\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t_{0}}V_{i}[s]\right]+\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s]]+O\left(\sqrt{\frac{\log t}{t}}\right).\end{split}$
Clearly, $\frac{1}{t}\sum_{s=1}^{t_{0}}V_{i}[s]\overset{\mathrm{a.s.}}{\to}0$, so by uniform integrability, $\mathbb{E}\left[\frac{1}{t}\sum_{s=1}^{t_{0}}V_{i}[s]\right]\to 0$. Therefore,
$\liminf_{t\to\infty}\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s]]\leq\liminf_{t\to\infty}\frac{1}{t}\sum_{s=1}^{t}\mathbb{E}[U_{i}[s]],$
proving the Nash equilibrium claim.
For the utility claim, let $\tilde{U}_{i}[s]$ denote the utility of agent $i$ in the imaginary game where no budget constraints are enforced and all agents $j$ are following a $\beta_{j}$-aggressive strategy. Observe that the $\tilde{U}_{i}[s]$ are i.i.d. $\mathrm{Bernoulli}(p_{i}v^{\star}(\beta_{i}))$. By the strong law of large numbers, $\frac{1}{t}\sum_{s=1}^{t}\tilde{U}_{i}[s]\overset{\mathrm{a.s}}{\to}p_{i}v^{\star}(\beta_{i})$. The utilities in the actual game satisfy
$\frac{1}{t}\sum_{s=1}^{t}U_{i}[s]=\frac{1}{t}\sum_{s=1}^{t_{0}}U_{i}[s]+\frac{1}{t}\sum_{s=t_{0}+1}^{t}U_{i}[s]=\frac{1}{t}\sum_{s=1}^{t_{0}}U_{i}[s]+\frac{1}{t}\sum_{s=t_{0}+1}^{t}\tilde{U}_{i}[s],$
which has the same limit as $t\to\infty$. ∎
D.4 Deferred Proof of Lemma 6.1
First, we note that the worst-case value distribution for robustness is a $\mathrm{Bernoulli}(\alpha_{i})$ value distribution.
Lemma D.6.
Assume player $i$ has a policy $\hat{\pi}_{i}$ such that if they had a $\hat{\mathcal{F}_{i}}=\mathrm{Bernoulli}(\alpha_{i})$ value distribution, regardless of the behavior of other agents $j\neq i$ they would obtain utility
$\frac{1}{T}\sum_{t=1}^{T}\hat{U}_{i}[t]\geq\lambda_{i}\hat{v}_{i}^{\star}$
with probability at least $1-O(1/T^{2})$ where $\hat{v}_{i}^{\star}=\alpha_{i}$ is the ideal utility of agent $i$ had they a $\mathrm{Bernoulli}(\alpha_{i})$ value distribution. Then, if, instead, player $i$ had an arbitrary value distribution $\mathcal{F}_{i}$, we can construct a policy $\pi_{i}$ such that regardless of the behavior of other agents $j\neq i$, player $i$ would obtain utility
$\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]\geq\lambda_{i}v_{i}^{\star}-O\left(\sqrt{\frac{\log T}{T}}\right)$
with probability at least $1-O(1/T^{2})$.
Proof.
Suppose agent $i$ has arbitrary value distribution $\mathcal{F}_{i}$. Construct the policy $\pi_{i}$ as follows. At each time $t$, agent $i$ will sample $\hat{V}_{i}[t]\sim\mathrm{Bernoulli}((\rho_{i}^{\alpha_{i}})^{\star}(V_{i}[t]))$, where $(\rho_{i}^{\alpha_{i}})^{\star}$ is agent $i$’s $\alpha_{i}$-ideal utility probability function with value distribution $\mathcal{F}_{i}$. (We define $\alpha_{i}$-ideal utility probability function in Definition A.1; informally, it is $1$ if $V_{i}[t]$ is in the top $\alpha_{i}$-quantile of $\mathcal{F}_{i}$ and 00 otherwise.) Then, agent $i$ will bid if and only if $\hat{\pi}_{i}$ would bid with the Bernoulli value $\hat{V}_{i}[t]$. In other words, the policy $\pi_{i}$ is simply following the policy $\hat{\pi}_{i}$ but with the Bernoulli values $\hat{V}_{i}[t]$ instead of the actual values.
Notice that the $\hat{V}_{i}[t]$ are indeed i.i.d. $\mathrm{Bernoulli}(\alpha_{i})$. By the hypothesis of the lemma,
$\frac{1}{T}\sum_{t=1}^{T}\hat{V}_{i}[t]\boldsymbol{1}\{i^{t}=i\}\geq\lambda_{i}\alpha_{i}$
on an event $E_{1}$ of probability at least $1-O(1/T^{2})$.
Let $\mathcal{H}_{t}$ denote the history up to time $t$. Let $\mathcal{G}_{t}$ be the $\sigma$-algebra generated by $\mathcal{H}_{t}$, $\hat{V}_{i}[t+1]$, and $i^{t+1}$. Define the $\mathcal{G}_{t}$-adapted process
$M[t]=\sum_{s=1}^{t}V_{i}[s]\boldsymbol{1}\{i^{s}=s\}-\frac{v_{i}^{\star}}{\alpha_{i}}\sum_{s=1}^{t}\hat{V}_{i}[s]\boldsymbol{1}\{i^{s}=s\}$
Observe that when agent $i$ uses the policy $\hat{\pi}_{i}$, everything in the mechanism is independent of the actual values $V_{i}[t]$ conditioned on the Bernoulli values $\hat{V}_{i}[t]$. Using this fact,
$\begin{split}\mathbb{E}[V_{i}[t]\boldsymbol{1}\{i^{t}=i\}\mid\mathcal{G}_{t-1}]&=\mathbb{E}[V_{i}[t]\mid\hat{V}_{i}[t]]\boldsymbol{1}\{i^{t}=i\}\\ &=\frac{\mathbb{E}[V_{i}[t]\boldsymbol{1}\{\hat{V}_{i}[t]=1\}]}{\Pr(\hat{V}_{i}[t]=1)}\hat{V}_{i}[t]\boldsymbol{1}\{i^{t}=i\}\\ &=\frac{\mathbb{E}[V_{i}[t](\rho_{i}^{\alpha_{i}})^{\star}(V_{i}[t])]}{\Pr(\hat{V}_{i}[t])=1)}\hat{V}_{i}[t]\boldsymbol{1}\{i^{t}=i\}\\ &=\frac{v_{i}^{\star}}{\alpha_{i}}\cdot\hat{V}_{i}[t]\boldsymbol{1}\{i^{t}=i\}.\end{split}$
Therefore, $M[t]$ is a $\mathcal{G}_{t}$-martingale. Let $\bar{v}$ be an upper bound on the distribution $\mathcal{F}_{i}$ (recall we assumed value distributions are bounded). By the Azuma-Hoeffding inequality, for any $\epsilon>0$,
$\Pr(M[T]\leq-\epsilon)\leq\exp\left(-\frac{2\epsilon^{2}}{\bar{v}^{2}T}\right).$
Set $\epsilon=\bar{v}\sqrt{T\ln T}$, so that the above is at most $O(1/T^{2})$. Let $E_{2}$ be the event that the above does not occur and let $E=E_{1}\cap E_{2}$. We have $\Pr(E)\geq 1-O(1/T^{2})$. On $E$,
$\frac{1}{T}\sum_{t=1}^{T}V_{i}[t]\boldsymbol{1}\{i^{t}=i\}=\frac{1}{T}M[T]+\frac{v_{i}^{\star}}{\alpha_{i}T}\sum_{t=1}^{T}\hat{V}_{i}[t]\boldsymbol{1}\{i^{t}=i\}\geq\lambda_{i}v_{i}^{\star}-O\left(\sqrt{\frac{\log T}{T}}\right),$
thereby proving the lemma. ∎
Now we can prove Lemma 6.1.
Proof of Lemma 6.1.
Let $\mathcal{H}_{t}$ denote the history up to time $t$. Let $\mathcal{G}_{t}$ be the $\sigma$-algebra generated by $\mathcal{H}_{t}$ and $b_{j}^{t+1}$ for $j\neq i$, i.e., make the bids of agents $j\neq i$ predictable processes with respect to $\mathcal{G}_{t}$. Define the $\mathcal{G}_{t}$-adapted process
$M_{i}^{t}=\sum_{s=1}^{t}\left(b_{i}^{s}\boldsymbol{1}\{i^{t}\neq i\}-\alpha_{i}\bar{p}\sum_{j\neq i}b_{j}^{s}\right).$
If $\sum_{j\neq i}b_{j}^{s}=0$, then
$\mathbb{E}\left[b_{i}^{s}\boldsymbol{1}\{i^{t}\neq i\}\,\middle|\,\mathcal{G}_{s-1}\right]=0.$
If $\sum_{j\neq i}b_{j}^{s}=1=b_{j}^{s}$, then
$\mathbb{E}\left[b_{i}^{s}\boldsymbol{1}\{i^{t}\neq i\}\,\middle|\,\mathcal{G}_{s-1}\right]\leq\alpha_{i}p_{j}^{\{i,j\}}\leq\alpha_{i}\bar{p}.$
If $\sum_{j\neq i}b_{j}^{s}\geq 2$, then
$\mathbb{E}\left[b_{i}^{s}\boldsymbol{1}\{i^{t}\neq i\}\,\middle|\,\mathcal{G}_{s-1}\right]\leq\alpha_{i}.$
In any case, since $\bar{p}\geq\frac{1}{2}$,
$\mathbb{E}\left[b_{i}^{s}\boldsymbol{1}\{i^{t}\neq i\}\,\middle|\,\mathcal{G}_{s-1}\right]\leq\alpha_{i}\bar{p}\sum_{j\neq i}b_{j}^{s}.$
Therefore, $M_{i}^{t}$ is a $\mathcal{G}_{t}$-supermartingale. Letting $\delta^{T}=\max_{j\neq i}\delta_{j}^{T}$, observe that
$\sum_{s=1}^{T}\sum_{j\neq i}b_{j}^{s}\leq(1-\alpha_{i})(1+\delta^{T})T\leq(1-\alpha_{i})T+O(\sqrt{T\log T}).$
So,
$\begin{split}\sum_{s=1}^{T}b_{i}^{s}\boldsymbol{1}\{i^{t}\neq i\}&\leq\sum_{s=1}^{T}\mathbb{E}[b_{i}^{s}\boldsymbol{1}\{i^{t}\neq i\}\mid\mathcal{G}_{s-1}]+M_{i}^{T}\\ &\leq\alpha_{i}\bar{p}\sum_{s=1}^{T}\sum_{j\neq i}b_{j}^{s}+M_{i}^{T}\\ &\leq\alpha_{i}\bar{p}(1-\alpha_{i})T+O(\sqrt{T\log T})+M_{i}^{T}.\end{split}$
This will be at most $\alpha_{i}\bar{p}(1-\alpha_{i})T+O(\sqrt{T\log T})$ on an event $E_{1}$ of probability at least $1-O\left(\frac{1}{T^{2}}\right)$ by Azuma-Hoeffding applied to $M_{i}^{T}$. Also, agent $i$ will bid at least $\alpha_{i}T-(\sqrt{T\log T})$ times on some event $E_{2}$ or probability at least $1-O\left(\frac{1}{T^{2}}\right)$ by standard Chernoff bounds. Let $E=E_{1}\cap E_{2}$. We lower bound agent $i$’s utility on $E$. Appealing to Lemma D.6, we assume without loss of generality that agent $i$ has a $\mathrm{Bernoulli}(\alpha_{i})$ value distribution. Then, the $\alpha_{i}$-aggressive strategy is simply the strategy of bidding when $V_{i}[t]=1$ if there is budget left. Agent $i$’s utility can then be bounded below on $E$ as
$\begin{split}\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]&=\frac{1}{T}\sum_{t=1}^{T}V_{i}[t]b_{i}^{t}\boldsymbol{1}\{i^{t}=i\}=\frac{1}{T}\sum_{t=1}^{T}b_{i}^{t}\boldsymbol{1}\{i^{t}=i\}\\ &=\frac{1}{T}\sum_{t=1}^{T}b_{i}^{t}-\frac{1}{T}\sum_{t=1}^{T}b_{i}^{t}\boldsymbol{1}\{i\neq i^{t}\}\geq\left(1-\alpha_{i}\bar{p}(1-\alpha_{i})\right)\alpha_{i}-O\left(\sqrt{\frac{\log T}{T}}\right).\end{split}$
Because the $\alpha_{i}$-ideal utility under a $\mathrm{Bernoulli}(\alpha_{i})$ value distribution is $\alpha_{i}$, this establishes the lemma. ∎
D.5 Completion of the Proof of Lemma 6.3
What is left to verify is that $f(\cdot,(y_{j})_{j\notin I})$ is Schur-concave for each $(y_{j})_{j\notin I}$ and $f((x_{i})_{i\in I},\cdot)$ is Schur-convex for each $(x_{i})_{i\in I}$. Notice that both $f(\cdot,(y_{j})_{j\notin I})$ and $f((x_{i})_{i\in I},\cdot)$ are symmetric functions. By the Schur–Ostrowski criterion, it suffices to show that
$(x_{i_{1}}-x_{i_{2}})\left(\frac{\partial f}{\partial x_{i_{1}}}-\frac{\partial f}{\partial x_{i_{2}}}\right)\leq 0$
and
$(y_{j_{1}}-y_{j_{2}})\left(\frac{\partial f}{\partial y_{j_{1}}}-\frac{\partial f}{\partial y_{j_{2}}}\right)\geq 0$
for all $((x_{i})_{i\in I},(y_{j})_{j\notin I})\in K$.
First, let us prove the following lemma.
Lemma D.7.
Let $\phi:[0,b]\to[0,\infty]$ be convex and nondecreasing with $\phi(0)=0$. For any nonnegative $x_{1},\dots,x_{n}$ with $\sum_{i=1}^{n}x_{i}\leq b$,
$\sum_{i=1}^{n}\phi(x_{i})\leq\phi\left(\sum_{i=1}^{n}x_{i}\right).$
Proof.
Define $\bar{x}=\sum_{i=1}^{n}x_{i}$, and consider the optimization problem below.
$\max_{(y_{i})}$ $\sum_{i=1}^{n}\phi(y_{i})$
s.t. $\sum_{i=1}^{n}y_{i}\leq\bar{x}$
$y_{i}\geq 0$ $\forall i\in[n]$
Since $\phi$ is convex, an optimal solution $(y_{i}^{\star})$ lies on an extreme point of the feasible region. Since the feasible region is a polytope defined by $n+1$ constraints, at least $n$ of them must be tight at an extreme point. Since $\phi$ is nondecreasing, this implies that there exists a unique $i^{\star}$ such that $y^{\star}_{i^{\star}}=\bar{x}$ and all other $y^{\star}_{i}=0$. Notice that $(x_{i})$ is a feasible solution to the optimization problem. It follows that
$\sum_{i=1}^{n}\phi(x_{i})\leq\sum_{i=1}^{n}\phi(y^{\star}_{i})=\phi(y^{\star}_{i^{\star}})+\sum_{i\neq i^{\star}}\phi(y^{\star}_{i})=\phi(\bar{x})=\phi\left(\sum_{i=1}^{n}x_{i}\right).$
To verify that $f(\cdot,(y_{j})_{j\notin I})$ is Schur-concave for each $(y_{j})_{j\notin I}$, compute
$\begin{split}(x_{i_{1}}-x_{i_{2}})\left(\frac{\partial f}{\partial x_{i_{1}}}-\frac{\partial f}{\partial x_{i_{2}}}\right)=\frac{1}{2}(x_{i_{1}}-x_{i_{2}})^{2}\prod_{i\in I\setminus\{i_{1},i_{2}\}}(1-x_{i})\left(-2+2\prod_{j\notin I}(1-y_{j})\right.\\ \left.+\prod_{j\notin I}(1-y_{j})\sum_{i\in I\setminus\{i_{1},i_{2}\}}\frac{x_{i}}{1-x_{i}}(-1+X)\right)\leq 0\end{split}$
where the inequality comes from the inequalities $-2+2\prod_{j\notin I}(1-y_{j})\leq 0$ and $-1+X\leq 0$.
To verify that $f((x_{i})_{i\in I},\cdot)$ is Schur-convex for each $(x_{i})_{i\in I}$, compute
$\begin{split}(y_{j_{1}}-y_{j_{2}})\left(\frac{\partial f}{\partial y_{j_{1}}}-\frac{\partial f}{\partial y_{j_{2}}}\right)=\frac{1}{2}\prod_{i\in I}(1-x_{i})\prod_{j\notin I\cup\{j_{1},j_{2}\}}(1-y_{j})\left(2X-\sum_{i\in I}\frac{x_{i}}{1-x_{i}}(1-X)\right).\end{split}$ (29)
By Lemma D.7 applied to the function $x\mapsto\frac{x}{1-x}$ on $[0,1]$,
$2X-\sum_{i\in I}\frac{x_{i}}{1-x_{i}}(1-X)\geq 2X-\frac{\sum_{i\in I}x_{i}}{1-\sum_{i\in I}x_{i}}(1-X)=2X-\frac{X}{1-X}(1-X)=X\geq 0$
Therefore, (29) is nonnegative so $f((x_{i})_{i\in I},\cdot)$ is Schur-convex.
D.6 Deferred Proofs of Propositions 6.4 and 6.5
Notice that Proposition 6.4 is a special case of Lemma 6.5, so we only need to prove Lemma 6.5.
Proof of Lemma 6.5.
The other agents $j\neq i$ will use the following strategy. At each time $t$, if every $j\neq i$ has budget remaining, either no $j\neq i$ will bid or a single agent $j$ will bid, where an agent $j$ bids with probability $\alpha_{j}$ (and therefore no agent will bid with probability $1-\sum_{j\neq i}\alpha_{j}=\alpha_{i}$). Their strategy will be independent across times, but notice that the agents’ bidding are very much not independent across agents.
For any single agent $j$, their bids are i.i.d. $\mathrm{Bernoulli}(\alpha_{j})$ across time conditioned on them having budget remaining. Let $E$ be the event that no agent $j\neq i$ runs out of budget. For the same reason as in Lemma D.2, the probability that agent $j$ runs out of budget is at most $O(1/T^{2})$, so $\Pr(E)\geq 1-O(1/T^{2})$. Using the idea and notation of the imaginary game where budgets are not enforced as introduced in Section D.2, agent $i$’s expected utility can be bounded as
$\begin{split}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]\leq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]\boldsymbol{1}_{E^{c}}]\leq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]+O\left(\frac{1}{T^{2}}\right).\end{split}$
Thus, we just need to bound the imaginary utility $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]$.
By the strategy of the other agents, (we use the notation for $\beta$-ideal utility as in Definition A.1)
$\begin{split}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]&=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]b_{i}^{t}\boldsymbol{1}\{i^{t}=i\}]=\frac{1}{T}\sum_{t=1}^{T}\left(\sum_{j\neq i}\alpha_{j}p_{j}^{\{i,j\}}+\alpha_{i}\right)\mathbb{E}[V_{i}[t]b_{i}^{t}]\\ &\leq\left(\sum_{j\neq i}\alpha_{j}p_{j}^{\{i,j\}}+\alpha_{i}\right)v_{i}^{\star}((1+\delta_{i}^{T})\alpha_{i})\\ &\leq\left(\sum_{j\neq i}\alpha_{j}p_{j}^{\{i,j\}}+\alpha_{i}\right)v_{i}^{\star}(\alpha_{i})+O\left(\sqrt{\frac{\log T}{T}}\right),\end{split}$
using Lemma D.1 for the first inequality and Lemma D.4 for the second. The above and (D.6) imply the lemma statement. ∎
D.7 Deferred Proof of Theorem 6.6
Proof of Theorem 6.6.
First, compute
$\begin{split}\sum_{i=1}^{n}\sum_{j\neq i}\alpha_{i}\alpha_{j}p_{i}^{\{i,j\}}=\sum_{i=1}^{n}\sum_{j\neq i}\alpha_{i}\alpha_{j}(1-p_{j}^{\{i,j\}})=\sum_{i=1}^{n}\sum_{j\neq i}\alpha_{i}\alpha_{j}-\sum_{i=1}^{n}\sum_{j\neq i}\alpha_{i}\alpha_{j}p_{j}^{\{i,j\}}\\ =1-\sum_{i=1}^{n}\alpha_{i}^{2}-\sum_{i=1}^{n}\sum_{j\neq i}\alpha_{i}\alpha_{j}p_{i}^{\{i,j\}}.\end{split}$
Solving for $\sum_{i=1}^{n}\sum_{j\neq i}\alpha_{i}\alpha_{j}p_{i}^{\{i,j\}}$,
$\sum_{i=1}^{n}\sum_{j\neq i}\alpha_{i}\alpha_{j}p_{i}^{\{i,j\}}=\frac{1}{2}-\frac{1}{2}\sum_{i=1}^{n}\alpha_{i}^{2}.$
Using the above and Lemma 6.5,
$\begin{split}\lambda=\lambda\sum_{i=1}^{n}\alpha_{i}&\leq\sum_{i=1}^{n}\left(\alpha_{i}+\sum_{j\neq i}\alpha_{j}p_{i}^{\{i,j\}}\right)\alpha_{i}+O\left(\sqrt{\frac{\log T}{T}}\right)\\ &=\sum_{i=1}^{n}\alpha_{i}^{2}+\left(\frac{1}{2}-\frac{1}{2}\sum_{i=1}^{n}\alpha_{i}^{2}\right)+O\left(\sqrt{\frac{\log T}{T}}\right)\\ &=\frac{1}{2}+\frac{1}{2}\sum_{i=1}^{n}\alpha_{i}^{2}+O\left(\sqrt{\frac{\log T}{T}}\right).\end{split}$
D.8 Completion of the Proof in Section C.2
Let $B$ be the set of bid rates and interim allocation probabilities $(\beta_{i},p_{i})$ that satisfy Border’s Criterion. In Section C.2, we showed that the set of $(\beta_{i},p_{i})\in B$ satisfying (9) with strict inequality for $I\subsetneq[n]$ are induced by allocation probabilities. We now give a routine topological argument to show that we can extend this proof to all of $B$.
The set of $(\beta_{i},p_{i})$ that are induced by allocation probabilities is closed (as a subset of Euclidean space): the set
$\left\{(\beta_{i},p_{i},p_{i}^{S}):p_{i}=\sum_{S\subseteq[n]:i\in S}p_{i}^{S}\left(\prod_{j\in S\setminus\{i\}}\beta_{j}\right)\left(\prod_{j\in[n]\setminus S}(1-\beta_{j})\right)\right\}$
is compact being the inverse image of a closed set under a continuous mapping, and the set of $(\beta_{i},p_{i})$ that are induced by allocation probabilities is just the projection of this set onto the first two coordinates. Hence, it suffices to show that the set of $(\beta_{i},p_{i})\in B$ satisfying (9) with strict inequality for $I\subseteq[n]$ is dense in $B$. Given any such $(\beta_{i},p_{i})\in B$, letting $\epsilon>0$, define $(\beta_{i}^{\prime},p_{i}^{\prime})$ by
$\beta_{i}^{\prime}$ $=\beta_{i}+\epsilon$
$p_{i}^{\prime}$ $=\frac{\beta_{i}p_{i}}{\beta_{i}^{\prime}}\cdot\frac{\prod_{j=1}^{n}(1-\beta_{j}^{\prime})}{\prod_{j=1}^{n}(1-\beta_{j})}.$
Then, $(\beta_{i}^{\prime},p_{i}^{\prime})$ can be made arbitrarily close to $(\beta_{i},p_{i})$. They can be seen to satisfy Border’s Criterion where (9) is strict for $I\subsetneq[n]$ as follows. We have increased $\beta_{i}$ and decreased $p_{i}$ to get $(\beta_{i}^{\prime},p_{i}^{\prime})$ in such a way that both the left-hand side and right-hand side of (8) increase in the same amount to maintain equality. In (9) for $I\subsetneq[n]$, considering the changes caused by each coordinate $i\in I$ one at a time, the left-hand side increases by the same amount as in (8) but the right-hand side increases by a larger amount since the partial derivative of the right-hand side with respect to some $\beta_{i}\in I$ is (strictly) decreasing in $I$ (when all $\beta_{j}>0$ for $j\in I$). Therefore, with the $(\beta_{i}^{\prime},p_{i}^{\prime})$, (9) holds with strict inequality.